### v1

# On the complementation of spaces of $\mathcal{I}$-null sequences

Abstract.
We study the complementation (in $\ell_{\infty}$) of the Banach space $c_{0,\mathcal{I}}$, consisting of all bounded sequences $(x_{n})$ that $\mathcal{I}$-converge to 00, endowed with the supremum norm, where $\mathcal{I}$ is an ideal of subsets of $\mathbb{N}$. We show that the complementation of these spaces is related to a condition requiring that the ideal is the intersection of a countable family of maximal ideals, which we refer to as $\omega$-maximal ideals. We prove that if $c_{0,\mathcal{I}}$ admits a projection satisfying a certain condition, then $\mathcal{I}$ must be a special type of $\omega$-maximal ideal. Additionally, we characterize when the quotient space $c_{0,\mathcal{J}}/c_{0,\mathcal{I}}$ is finite-dimensional for two ideals $\mathcal{I}\subsetneq\mathcal{J}$.
1. Introduction
An ideal on $\mathbb{N}$ is a collection $\mathcal{I}$ of subsets of $\mathbb{N}$ closed under finite unions and taking subsets of its elements. A sequence $(x_{n})$ in a Banach space $X$ is said to be $\mathcal{I}$-convergent to $x\in X$, denoted as $\mathcal{I}$-$\lim x_{n}=x$, if for each $\varepsilon>0$, the set $\{n\in\mathbb{N}\,\colon\,\|x_{n}-x\|\geq\varepsilon\}$ belongs to $\mathcal{I}$. When $\mathcal{I}$ is ${\sf Fin}$, the ideal of finite subsets of $\mathbb{N}$, we have the classical convergence in $X$. For this reason, it is natural—and we will adopt this assumption—to require that ${\sf Fin}$ is contained in every ideal under consideration. The $\mathcal{I}$-convergence was introduced in [16], although many authors had already studied this concept in particular cases and in different contexts (see, for instance, [2, 3, 9, 10, 17]). We are interested in the following space
$c_{0,\mathcal{I}}=\{(x_{n})\in\ell_{\infty}\,\colon\,\mathcal{I}-\lim x_{n}=0\}.$
We showed in [23] that $c_{0,\mathcal{I}}$ is a closed subspace of $\ell_{\infty}$, and that some of its Banach and Banach-lattice properties are closely related to the combinatorial and topological properties of the ideal $\mathcal{I}$. For instance, a closed sublattice of $\ell_{\infty}$ is an ideal exactly when it is of the form $c_{0,\mathcal{I}}$ for some ideal $\mathcal{I}$ on $\mathbb{N}$. Furthermore, $c_{0,\mathcal{I}}$ and $c_{0,\mathcal{J}}$ are isometric if, and only if, $\mathcal{I}$ and $\mathcal{J}$ are isomorphic. The main objective of this paper is to investigate the phenomenon of complementation of $c_{0,\mathcal{I}}$ in $\ell_{\infty}$. Some results in this direction are already known.
We call a proper ideal $\mathcal{I}$ complemented if $c_{0,\mathcal{I}}$ is complemented in $\ell_{\infty}$. Leonetti [18] proved that any meager ideal is not complemented, where a meager ideal refers to an ideal that is meager as a subset of the Cantor cube $\{0,1\}^{\mathbb{N}}$, identified via characteristic functions. The key element of his argument is the existence of uncountable families of subsets in $\mathcal{P}(\mathbb{N})\setminus\mathcal{I}$ such $A\cap B\in\mathcal{I}$ for any $A,B$ in the family (the so-called $\mathcal{I}$-AD families). We show that if $\mathcal{I}$ is complemented, then any $\mathcal{I}$-AD family is at most countable and therefore $\mathcal{I}$ is not meager and does not have the Baire property as a subset of $\{0,1\}^{\mathbb{N}}$.
On the other hand, Kania [14] observed that the intersection of a finite collection of maximal ideals is complemented (for a proof, see [23]). We call an ideal $\omega$-maximal if it can be written as a countable intersection of maximal ideals. We extend Kania’s result to certain special $\omega$-maximal ideals.
We say that an ideal is strongly $\omega$-maximal if there exists a collection $\{\mathcal{I}_{n}\colon n\in\mathbb{N}\}$ of maximal ideals such that $\mathcal{I}=\bigcap_{n}\mathcal{I}_{n}$, and the family $\{\mathcal{I}_{n}^{*}\colon n\in\mathbb{N}\}$ is discrete in $\beta\mathbb{N}$ (where $\mathcal{I}^{*}$ denotes the dual filter and $\beta\mathbb{N}$ is the Stone–Čech compactification of $\mathbb{N}$). We provide an example of an $\omega$-maximal ideal that is not strongly $\omega$-maximal. Since $\omega$-maximal ideals play a central role in our results, we present additional properties about them in Section 3. The dual notion of a filter represented as an intersection of a finite or countable family of ultrafilters has been recently studied by Bergman [4] and Kadets, Seliutin and Tryba [15]. In particular, we show that the notion introduced in [15] of a filter admitting a minimal countable representation corresponds to our notion of strongly $\omega$-maximal ideal (see Remark 3.6).
In Section 4, we study the quotient $c_{0,\mathcal{J}}/c_{0,\mathcal{I}}$ when $\mathcal{I}\subsetneq\mathcal{J}$ are ideals. We provide a combinatorial characterization of the finite-dimensionality of $c_{0,\mathcal{J}}/c_{0,\mathcal{I}}$. In particular, we obtain that $\ell_{\infty}/c_{0,\mathcal{I}}$ is finite-dimensional if and only if $\mathcal{I}$ is a finite intersection of maximal ideals.
In the last section, we discuss the problem of the complementation of $c_{0,\mathcal{I}}$ in $\ell_{\infty}$. One of our main results is that any strongly $\omega$-maximal ideal is complemented (see Theorem 5.8). We provide two different proofs of this result. The first one (see Theorem 5.2) is shorter but considerably less informative, as it relies on a classical theorem of Lindenstrauss stating that a closed subspace of $\ell_{\infty}$ is complemented if and only if it is isomorphic to $\ell_{\infty}$. Furthermore, we characterize strongly $\omega$-maximal ideals in terms of the type of projections on their associated space $c_{0,\mathcal{I}}$ (see Theorem 5.20).
In addition, we show that if $c_{0,\mathcal{I}}$ is complemented in $\ell_{\infty}$ by a projection satisfying an extra condition, then $\mathcal{I}$ must be an $\omega$-maximal ideal (see Theorem 5.18). Whether the converse holds—that is, whether every $\omega$-maximal ideal is complemented—remains an open question.
We show that $c_{0}$ is not complemented in $c_{0,\mathcal{J}}$ for any ideal $\mathcal{J}$ properly extending $\mathsf{Fin}$. This property is shared by all ideals $\mathcal{I}$ such that $\mathcal{I}\restriction A$ is Baire measurable on $2^{A}$ for any $A\notin\mathcal{I}$. For instance, all analytic ideals have this hereditary property.
Finally, we present several examples of ideals which are not complemented.
2. Preliminaries
We will use standard terminology and notation for Banach lattices and Banach space theory. For unexplained definitions and notations, we refer to [1, 24]. The scalar field is denoted by $\mathbb{K}$. All Banach lattices analyzed here are assumed to be real. However, our results can be extended to complex Banach lattices in the usual manner [24, Chapter 2, p. 133]. If $X$ and $Y$ are isomorphic Banach spaces, we write $X\sim Y$. If $E$ is a closed subspace of a Banach space $X$, we say that $E$ is complemented in $X$ if there is a continuous onto operator $P\colon X\to E$ such that $P^{2}=P$, or equivalently, there is a closed subspace $W$ of $X$ such that $X=E\oplus W$. In addition, if $E$ and $X$ are Banach lattices, $P$ is called positive if $Px\geq{\bf 0}$ for all $x\geq{\bf 0}$.
An ideal $\mathcal{I}$ on a set $X$ is a collection of subsets of $X$ satisfying:
  1. (1)
    $\emptyset\in\mathcal{I}$;
  2. (2)
    If $A\subseteq B$ and $B\in\mathcal{I}$, then $A\in\mathcal{I}$;
  3. (3)
    If $A,B\in\mathcal{I}$, then $A\cup B\in\mathcal{I}$.
We always assume that every finite subset of $X$ belongs to $\mathcal{I}$. The dual filter of an ideal $\mathcal{I}$ is denoted by $\mathcal{I}^{*}$ and consists of all sets of the form $X\setminus A$ for some $A\in\mathcal{I}$. The co-ideal $\mathcal{I}^{+}$ is the collection $\mathcal{P}(X)\setminus\mathcal{I}$.
When $X$ is countable, an ideal $\mathcal{I}$ can be conveniently seen as a subset of the Cantor cube $\{0,1\}^{X}$ with the compact metric topology. This allows us to consider when $\mathcal{I}$ is a meager subset of the Cantor cube. The following is a very useful result:
Theorem 2.1 (Jalali-Naini, Talagrand [13, 26]).
Let $\mathcal{I}$ be a proper ideal on $\mathbb{N}$. The following statements are equivalent:
  1. (1)
    $\mathcal{I}$ is meager.
  2. (2)
    $\mathcal{I}$ has the Baire property.
  3. (3)
    There is a partition $\{F_{k}\,\colon\,k\in\mathbb{N}\}$ of $\mathbb{N}$ into finite sets such that for every $M\subseteq\mathbb{N}$ infinite we have $\bigcup_{k\in M}F_{k}\not\in\mathcal{I}$.
An ideal $\mathcal{I}$ is maximal if $\mathcal{P}(X)$ is the only ideal properly extending $\mathcal{I}$; equivalently, if $\mathcal{I}^{*}$ is an ultrafilter. Notice that $\mathcal{I}$ is maximal if $\mathcal{I}^{*}=\mathcal{I}^{+}$. For $A\subseteq X$, we denote the restriction of $\mathcal{I}$ to $A$ by $\mathcal{I}\restriction A=\{A\cap B\,\colon B\in\mathcal{I}\}$ which is an ideal on $A$. Let $\mathcal{A}$ and $\mathcal{B}$ families of sets, we denote by $\mathcal{A}\sqcup\mathcal{B}$ the collection $\{A\cup B\,\colon\,A\in\mathcal{A},B\in\mathcal{B}\}.$ An $\mathcal{I}$-AD family is a collection $\mathcal{A}\subseteq\mathcal{I}^{+}$ such that $A\cap B\in\mathcal{I}$ for every two different sets $A,B\in\mathcal{A}$. Two ideals $\mathcal{I}$ and $\mathcal{J}$ on $X$ and $Y$, respectively, are isomorphic, if there is a bijection $f:X\to Y$ such that $f[E]\in\mathcal{J}$ for all $E\in\mathcal{I}$.
The following observations will be needed in the sequel, its proof is straightforward.
Lemma 2.2.
Let $\mathcal{I}$ be an ideal on $\mathbb{N}$ and $A\in\mathcal{I}^{+}$.
  1. (1)
    If $\mathcal{I}\restriction A$ is a maximal ideal on $A$, then $\mathcal{J}=\mathcal{I}\restriction A\sqcup\mathcal{P}(A^{c})$ is maximal ideal on $\mathbb{N}$.
  2. (2)
    If $\mathcal{I}$ is maximal and $A\in\mathcal{I}^{+}$, then $\mathcal{I}=\mathcal{I}\restriction A\sqcup\mathcal{P}(A^{c})$.
  3. (3)
    The following assertions are equivalent:
    1. (a)
      $\mathcal{I}\restriction A$ is maximal on $A$;
    2. (b)
      Let $\mathcal{J}_{A}=\mathcal{I}\sqcup\mathcal{P}(A)$. Then $\mathcal{I}\restriction B$ is maximal on $B$ for all $B\in\mathcal{J}_{A}\setminus\mathcal{I}$.
Let $\{K_{n}:\;n\in F\}$ be a partition of a countable set $X$, where $F\subseteq\mathbb{N}$. For $n\in F$, let $\mathcal{I}_{n}$ be an ideal on $K_{n}$. The direct sum, denoted by $\bigoplus\limits_{n\in F}\mathcal{I}_{n}$, is defined as follows:
$A\in\bigoplus_{n\in F}\mathcal{I}_{n}\Leftrightarrow(\forall n\in F)(A\cap K_{n}\in\mathcal{I}_{n}).$
Notice that the direct sum $\bigoplus_{n}\mathcal{I}_{n}$ can also be naturally defined in $\mathbb{N}\times\mathbb{N}$. When $\mathcal{I}_{n}$ is isomorphic to $\mathcal{I}$ for all $n$, the direct sum is denoted by $\mathcal{I}^{\omega}$.
If $\mathcal{I}$ is a maximal ideal on $\mathbb{N}$ and $K\not\in\mathcal{I}$, it is easy to see that $A\in\mathcal{I}$ if and only if $A\cap K\in\mathcal{I}$. From this, we have the following observation that will be used later on.
Lemma 2.3.
Let $\{\mathcal{I}_{n}\,\colon\,n\in F\}$ be a countable collection of maximal ideals on $\mathbb{N}$ and $\{K_{n}\,\colon\,n\in F\}$ be a family of pairwise disjoint subsets of $\mathbb{N}$ such that $K_{n}\not\in\mathcal{I}_{n}$ for each $n\in F$. Then, $A\in\bigcap_{n\in F}\mathcal{I}_{n}$ if and only if $A\cap K_{n}\in\mathcal{I}_{n}$ for every $n\in F$. In particular, if $\{K_{n}\,\colon\,n\in F\}$ is a partition of $\mathbb{N}$ such that $K_{n}\not\in\mathcal{I}_{n}$ for each $n\in F$, then
$\bigoplus_{n\in F}(\mathcal{I}_{n}\restriction K_{n})\,=\,\bigcap_{n\in F}\mathcal{I}_{n}.$
Recall that $\beta\mathbb{N}$ is the Stone-Čech compactification of $\mathbb{N}$ which is usually identified with the collection of all ultrafilters on $\mathbb{N}$. For a set $A\subseteq\mathbb{N}$, we let $A^{*}=\{p\in\beta\mathbb{N}\,\colon\,A\in p\}$. The family $\{A^{*}\,\colon\,A\subseteq\mathbb{N}\}$ defines a basis for the topology of $\beta\mathbb{N}$. As usual, we identify each $n\in\mathbb{N}$ with the principal ultrafilter $\{A\subseteq\mathbb{N}:n\in A\}$. Notice that every principal ultrafilter is an isolated point of $\beta\mathbb{N}$.
If ${\bf x}=(x_{n})\in\ell_{\infty}$ and $\varepsilon>0$, we will use throughout the whole paper the following notation:
$A(\varepsilon,{\bf x})=\{n\in\mathbb{N}\,\colon\,|x_{n}|\geq\varepsilon\}.$
Notice that ${\bf x}\in c_{0,\mathcal{I}}$ if and only if $A(\varepsilon,{\bf x})\in\mathcal{I}$ for all $\varepsilon>0$.
3. $\omega$-maximal ideals
Every ideal on $\mathbb{N}$ is easily seen to be equal to an intersection of a collection of $2^{\aleph_{0}}$ many maximal ideals. Therefore, an ideal $\mathcal{I}$ on $\mathbb{N}$ is called $\kappa$-maximal, for $\kappa$ a cardinal, if $\mathcal{I}=\bigcap_{\alpha<\kappa}\mathcal{I}_{\alpha}$ for some maximal ideals $\mathcal{I}_{\alpha}$, for $\alpha<\kappa$ and $\kappa$ has the smallest possible value. Clearly, the interesting cases are the $\kappa$-maximal ideals with $\kappa<2^{\aleph_{0}}$.
Since we are always assuming that ${\sf Fin}$ is contained in any ideal under consideration, every maximal ideal extending a given ideal is necessarily non-principal. It is worth keeping in mind that the intersection of less than $2^{\aleph_{0}}$ maximal ideals on $\mathbb{N}$ does not have the Baire property, and hence it is non-meager ([22], [26, Proposition 23]). Our main interest will be on $\omega$-maximal ideals since they are related to the complementation of $c_{0,\mathcal{I}}$ in $\ell_{\infty}$.
We say that a maximal ideal $\mathcal{J}$ is a limit point of a set $\mathfrak{D}$ of maximal ideals, if $\mathcal{J}^{*}$ is a limit point of $\mathfrak{D}^{*}\coloneqq\{\mathcal{I}^{*}:\mathcal{I}\in\mathfrak{D}\}$ in $\beta\mathbb{N}$. We say that a countable collection of maximal ideals $\{\mathcal{I}_{n}:n\in\mathbb{N}\}$ is discrete if $\{\mathcal{I}^{*}_{n}:n\in\mathbb{N}\}$ is a discrete subset of $\beta\mathbb{N}$. An ideal $\mathcal{I}$ is strongly $\omega$-maximal if there is a discrete collection $\{\mathcal{I}_{n}:n\in\mathbb{N}\}$ of maximal ideals on $\mathbb{N}$ such that $\mathcal{I}=\bigcap_{n}\mathcal{I}_{n}$. In this section, all discrete collections of maximal ideals are assumed to be infinite.
Now, we present a useful characterization of strongly $\omega$-maximal ideals. Since every finite subset of $\beta\mathbb{N}$ is discrete, the following result also applies to $k$-maximal ideals for any positive integer $k$, a case that was already shown by A. Millán [21] and Bergman [4]. Millán’s work was particularly helpful in understanding certain properties of $\omega$-maximal ideals.
Proposition 3.1.
Let $\{\mathcal{I}_{k}\,\colon\,k\in\mathbb{N}\}$ be a collection of maximal ideals on $\mathbb{N}$. Then, $\{\mathcal{I}_{k}\,\colon\,k\in\mathbb{N}\}$ is discrete if and only if there is a partition $\{A_{k}\,\colon\,k\in\mathbb{N}\}$ of $\mathbb{N}$ such that $A_{k}\in\mathcal{I}^{*}_{k}\cap(\bigcap_{j\in\mathbb{N}\setminus\{k\}}\mathcal{I}_{j})$ for all $k\in\mathbb{N}$.
Proof.
Suppose there is a partition $\{A_{k}\,\colon\,k\in\mathbb{N}\}$ of $\mathbb{N}$ such that $A_{k}\in\mathcal{I}^{*}_{k}\cap(\bigcap_{j\in\mathbb{N}\setminus\{k\}}\mathcal{I}_{j})$ for all $k\in\mathbb{N}$. Then $\{\mathcal{I}_{j}^{*}\,\colon\,j\in\mathbb{N}\}\cap A_{k}^{*}=\{\mathcal{I}_{k}^{*}\}$ for all $k\in\mathbb{N}$, thus $\{\mathcal{I}_{k}\,\colon\,k\in\mathbb{N}\}$ is discrete.
Now suppose that $\{\mathcal{I}_{j}\,\colon\,j\in\mathbb{N}\}$ is discrete. For each $k\in\mathbb{N}$, there exists $B_{k}\subseteq\mathbb{N}$ such that $\{\mathcal{I}_{j}^{*}\,\colon\,j\in\mathbb{N}\}\cap B_{k}^{*}=\{\mathcal{I}_{k}^{*}\}$, that is, $B_{k}\in\mathcal{I}_{k}^{*}\cap\mathcal{I}_{j}$ for all $j\in\mathbb{N}$ with $j\neq k$. Set $A_{1}=B_{1}$ and $A_{j}=B_{j}\setminus\bigcup_{1\leq i<j}B_{i}$ for all $j\geq 2$. Observe that if $k\in\mathbb{N}$, then $A_{k}\in\mathcal{I}_{j}$ for each $j\in\mathbb{N}$ with $j\neq k$. We claim that $A_{k}\in\mathcal{I}_{k}^{*}$ for all $k\in\mathbb{N}$. Indeed, if $k\in\mathbb{N}$ is given, we have $B_{k}=A_{k}\cup(\bigcup_{1\leq i<k}(B_{i}\cap B_{k}))$. Notice that $B_{i}\cap B_{k}\in\mathcal{I}_{k}$ for all $1\leq i<k$. Thus, $\bigcup_{1\leq i<k}(B_{i}\cap B_{k})\in\mathcal{I}_{k}$, and therefore $A_{k}\in\mathcal{I}_{k}^{*}$.
Finally, since $E=\mathbb{N}\setminus\bigcup_{j\in\mathbb{N}}A_{j}\in\mathcal{I}_{k}$ for all $k\in\mathbb{N}$, by substituting $B_{1}$ by $A_{1}=B_{1}\cup E$, we obtain that $\{A_{k}\,\colon\,k\in\mathbb{N}\}$ is a partition of $\mathbb{N}$ such that $A_{k}\in\mathcal{I}^{*}_{k}\cap(\bigcap_{j\in\mathbb{N}\setminus\{k\}}\mathcal{I}_{j})$ for each $k\in\mathbb{N}$. ∎
The following observation is crucial for what follows. Part (1) turned out to be known [12, Theorem 3.20] but we include a proof for sake of completeness.
Lemma 3.2.
Let $\mathfrak{D}$ be a collection of maximal ideals and $\mathcal{I}$ be a maximal ideal. Then
  1. (1)
    $\mathcal{I}^{*}\in\overline{\mathfrak{D}^{*}}$ if and only if $\bigcap\mathfrak{D}\subseteq\mathcal{I}$.
  2. (2)
[…]
Proof.
  1. (1)
    $\mathcal{I}^{*}\not\in\overline{\mathfrak{D}^{*}}$ if and only if there is $A\subseteq\mathbb{N}$ such that $A\in\mathcal{I}^{*}$ and $A^{*}\cap\mathfrak{D}^{*}=\emptyset$ if and only if there is $A\subseteq\mathbb{N}$ such that $A^{c}\not\in\mathcal{I}^{*}$ and $A^{c}\in\mathcal{J}^{*}$ for all $\mathcal{J}^{*}\in\mathfrak{D}^{*}$ if and only if $\bigcap\mathfrak{D}\not\subseteq\mathcal{I}$.
  2. (2)
    It follows from (1).∎
For each collection of maximal ideals $\mathfrak{E}=\{\mathcal{I}_{n}\,\colon\,n\in\mathbb{N}\}$ we define a family of infinite subsets of $\mathbb{N}$ which will be very helpfull for what follows:
$\mathcal{DI}(\mathfrak{E})=\{M\in[\mathbb{N}]^{\omega}\,\colon\,\mbox{$\{\mathcal{I}_{n}\,\colon\,n\in M\}$ is discrete}\}.$
For each $M\subseteq\mathbb{N}$, we let $\mathcal{I}_{M}=\bigcap_{n\in M}\mathcal{I}_{n}$.
Proposition 3.3.
Let $\mathfrak{E}=\{\mathcal{I}_{n}\,\colon\,n\in\mathbb{N}\}$ be a collection of maximal ideals on $\mathbb{N}$ and $A,M\subseteq\mathbb{N}$. Then
  1. (1)
    $\{\mathcal{I}_{n}^{*}\,\colon\,n\in A\}\subseteq\overline{\{\mathcal{I}_{m}^{*}\,\colon\,m\in M\}}$ if and only if $\mathcal{I}_{M}=\mathcal{I}_{M\cup A}$.
  2. (2)
    Let $M\in\mathcal{DI}(\mathfrak{E})$ and $n\in\mathbb{N}$. Then, $\mathcal{I}_{M}\not\subseteq\mathcal{I}_{n}$ if and only if $M\cup\{n\}\in\mathcal{DI}(\mathfrak{E})$.
  3. (3)
    Let $M\in\mathcal{DI}(\mathfrak{E})$. Then, $M$ is maximal in $\mathcal{DI}(\mathfrak{E})$ if and only if $\mathcal{I}_{\mathbb{N}}=\mathcal{I}_{M}$.
  4. (4)
    Let $M\subseteq\mathbb{N}$ and $D(M)=\{m\in M\,\colon\,\bigcap_{n\in M\setminus\{m\}}\mathcal{I}_{n}\not\subseteq\mathcal{I}_{m}\}$. Then $D(M)\in\mathcal{DI}(\mathfrak{E})$.
Proof.
(1) By Lemma 3.2 we have that
$\mathcal{I}_{M}=\mathcal{I}_{M\cup A}\Longleftrightarrow\mathcal{I}_{M}\subset\mathcal{I}_{A}\Longleftrightarrow\mathcal{I}_{n}^{*}\in\overline{\{{\mathcal{I}}_{m}^{*}\,\colon\,m\in M\}}\,\,\mbox{for all $n\in A$.}$
(2) By Lemma 3.2 we have
$\mathcal{I}_{M}\not\subseteq\mathcal{I}_{n}\Longleftrightarrow\mathcal{I}_{n}^{*}\not\in\overline{\{\mathcal{I}_{m}^{*}\,\colon\,m\in M\}}\Longleftrightarrow\mbox{$\{\mathcal{I}_{m}\,\colon\,m\in M\}\cup\{\mathcal{I}_{n}\}$ is discrete}\Longleftrightarrow M\cup\{n\}\in\mathcal{DI}(\mathfrak{E}).$
(3) Suppose $M$ is maximal in $\mathcal{DI}(\mathfrak{E})$. It suffices to show that $\mathcal{I}_{M}\subseteq\mathcal{I}_{\mathbb{N}}$, that is, $\mathcal{I}_{M}\subseteq\mathcal{I}_{n}$ for all $n\in\mathbb{N}$. Let $n\in\mathbb{N}$ be given which clearly can be assumed not in $M$. By the maximality of $M$, $M\cup\{n\}\not\in\mathcal{DI}(\mathfrak{E})$. Thus, by (2), $\mathcal{I}_{M}\subseteq\mathcal{I}_{n}$.
Conversely, suppose that $M$ is not maximal. Thus there is $n\in\mathbb{N}\setminus M$ such that $M\cup\{n\}\in\mathcal{DI}(\mathfrak{E})$. By (2) we have $\mathcal{I}_{M}\not\subseteq\mathcal{I}_{n}$. Thus $\mathcal{I}_{\mathbb{N}}\neq\mathcal{I}_{M}$.
(4) Notice that for all $m\in D(M)$ we have $\bigcap_{n\in D(M)\setminus\{m\}}\mathcal{I}_{n}\not\subseteq\mathcal{I}_{m}$. By Lemma 3.2 we conclude that $D(M)\in\mathcal{DI}(\mathfrak{E})$. ∎
Theorem 3.4.
Let $\mathfrak{E}=\{\mathcal{I}_{n}:n\in\mathbb{N}\}$ be a collection of maximal ideals on $\mathbb{N}$ and $\mathcal{I}=\bigcap_{n}\mathcal{I}_{n}$. Then, $\mathcal{I}$ is strongly $\omega$-maximal if and only if $\mathcal{DI}(\mathfrak{E})$ has a $\subseteq$-maximal element. Moreover, $\mathcal{DI}(\mathfrak{E})$ has at most one maximal element. In particular, a strongly $\omega$-maximal ideal admits a unique representation by a discrete collection of maximal ideals.
Proof.
If $M$ is a maximal element of $\mathcal{DI}(\mathfrak{E})$, by Proposition 3.3, $\mathcal{I}=\mathcal{I}_{M}$, thus $\mathcal{I}$ is strongly $\omega$-maximal.
Conversely, suppose $\mathcal{I}$ is strongly $\omega$-maximal and let $\mathfrak{D}=\{\mathcal{K}_{n}:n\in\mathbb{N}\}$ be a discrete collection of maximal ideals on $\mathbb{N}$ such that $\mathcal{I}=\bigcap_{n}\mathcal{K}_{n}$. By Proposition 3.1, there is a partition $\{A_{n}:n\in\mathbb{N}\}$ of $\mathbb{N}$ such that $A_{n}\not\in\mathcal{K}_{n}$ for all $n\in\mathbb{N}$. We will show that for each $n$ there is a unique $l_{n}$ such that $\mathcal{I}_{l_{n}}=\mathcal{K}_{n}$. Then, letting $M=\{l_{n}\,\colon\,n\in\mathbb{N}\}$ we have that $M\in\mathcal{DI}(\mathfrak{E})$, as $\mathfrak{D}$ is discrete, $\mathcal{I}=\mathcal{I}_{M}$ and $M$ is maximal by Proposition 3.3.
Given $n\in\mathbb{N}$, as $A_{n}\not\in\mathcal{I}$ for all $n$, there is $m$ such that $A_{n}\not\in\mathcal{I}_{m}$. We claim that $\mathcal{I}_{m}\restriction A_{n}=\mathcal{K}_{n}\restriction A_{n}$. Indeed, as the $A_{n}$’s are pairwise disjoint, $\mathcal{I}\restriction A_{n}=\mathcal{K}_{n}\restriction A_{n}$. In particular, $\mathcal{K}_{n}\restriction A_{n}\subseteq\mathcal{I}_{m}\restriction A_{n}$, but $\mathcal{I}_{m}\restriction A_{n}$ and $\mathcal{K}_{m}\restriction A_{n}$ are maximal ideals on $A_{n}$, thus $\mathcal{I}_{m}\restriction A_{n}=\mathcal{K}_{n}\restriction A_{n}$. Therefore $\mathcal{I}_{m}=\mathcal{K}_{n}$. Notice that this argument shows that for all $n$ there is a unique $l_{n}$ such that $\mathcal{I}_{l_{n}}=\mathcal{K}_{n}$.
Finally, suppose $M_{1}$ and $M_{2}$ are two maximal elements of $\mathcal{DI}(\mathfrak{E})$. By Proposition 3.3, $\mathcal{I}=\mathcal{I}_{M_{1}}=\mathcal{I}_{M_{2}}$. Let $\{A_{n}:n\in M_{1}\}$ be a partition of $\mathbb{N}$ such that $A_{n}\not\in\mathcal{I}_{n}$ for all $n\in M_{1}$. Let $n_{1}\in M_{1}\setminus M_{2}$. We have shown above that $n_{1}$ is the unique $m\in\mathbb{N}$ such that $A_{n_{1}}\not\in\mathcal{I}_{m}$. Thus $A_{n_{1}}\in\mathcal{I}_{m}$ for all $m\in M_{2}$. Thus $A_{n_{1}}\in\mathcal{I}_{M_{2}}=\mathcal{I}$, which contradicts that $A_{n_{1}}\not\in\mathcal{I}_{n_{1}}$.
For the last claim, suppose $\mathfrak{E}_{1}$ and $\mathfrak{E}_{2}$ are two discrete countable collections of maximal ideals such that $\bigcap\mathfrak{E}_{1}=\bigcap\mathfrak{E}_{2}$. Let $\{J_{n}:n\in\mathbb{N}\}$ be an enumeration (without repetitions) of $\mathfrak{E}_{1}\cup\mathfrak{E}_{2}$. Since $\bigcap\mathfrak{E}_{1}=\bigcap(\mathfrak{E}_{1}\cup\mathfrak{E}_{2})$, $\mathcal{DI}(\mathfrak{E}_{1}\cup\mathfrak{E}_{2})$ has a unique maximal element $M$. Suppose there is $n_{0}$ such that $J_{n_{0}}\in\mathfrak{E}_{2}\setminus\mathfrak{E}_{1}$. Let $L=\{n\in\mathbb{N}:J_{n}\in\mathfrak{E}_{2}\}$ and notice that $L\in\mathcal{DI}(\mathfrak{E}_{1}\cup\mathfrak{E}_{2})$ is maximal by Lemma 3.3 and thus $n_{0}\in M$. On the other hand, $\bigcap\{J_{n}:n\in M\setminus\{n_{0}\}\}=\bigcap\mathfrak{E}_{2}\subseteq J_{n_{0}}$ and, by Lemma 3.2, $M$ is not discrete, a contradiction. ∎
We recall that a topological space $X$ is scattered if every non-empty subspace of $X$ has an isolated point. We say that a countable collection $\mathfrak{E}$ of maximal ideals on $\mathbb{N}$ is scattered if $\mathfrak{E}^{*}$ is a scattered subspace of $\beta\mathbb{N}$.
Theorem 3.5.
Let $\mathcal{I}$ be an ideal on $\mathbb{N}$. Then $\mathcal{I}=\bigcap\mathfrak{E}$ for a countable scattered collection $\mathfrak{E}$ of maximal ideals on $\mathbb{N}$ if and only if $\mathcal{I}$ is strongly $\omega$-maximal.
Proof.
Suppose $\mathfrak{E}^{*}$ is scattered. Since the collection of isolated points of $\mathfrak{E}^{*}$ is discrete and dense in $\mathfrak{E}^{*}$ (see [25, p. 150]), there is $M\in\mathcal{DI}(\mathfrak{E})$ such that $\{\mathcal{I}^{*}:\,n\in M\}$ is dense in $\mathfrak{E}^{*}$, then easily $M$ is maximal in $\mathcal{DI}(\mathfrak{E})$. Thus by Theorem 3.4, $\mathcal{I}$ is strongly $\omega$-maximal. The converse is obvious. ∎
Remark 3.6.
[…]
We present two examples. The first one shows that a strongly $\omega$-maximal ideal can also be represented by a non-discrete countable collection of maximal ideals. The second one provides an example of a $\omega$-maximal ideal which is not strongly $\omega$-maximal.
Example 3.7.
Let $\mathfrak{E}=\{\mathcal{I}_{n}:n\in\mathbb{N}\}$ be a discrete collection of maximal ideals on $\mathbb{N}$ and $\mathcal{I}=\bigcap_{n}\mathcal{I}_{n}$. Let $\{A_{n}:n\in\mathbb{N}\}$ be a partition of $\mathbb{N}$ such that $A_{n}\not\in\mathcal{I}_{n}$ for all $n\in\mathbb{N}$. Let $\mathcal{J}_{0}$ be the ideal generated by $\mathcal{I}\cup\{A_{n}:n\in\mathbb{N}\}$. Then $\mathcal{J}_{0}$ is non-trivial, in fact, if $\mathbb{N}=B\cup A_{0}\cup\ldots\cup A_{m}$, then $A_{m+1}\subseteq B$, thus $B\not\in\mathcal{I}$ and hence $\mathbb{N}\not\in\mathcal{J}_{0}$. Let $\mathcal{J}$ be a maximal ideal extending $\mathcal{J}_{0}$. Then $\mathcal{I}=\bigcap(\mathfrak{E}\cup\{\mathcal{J}\})$ but $\mathfrak{E}\cup\{\mathcal{J}\}$ is not discrete.
Example 3.8.
Let $\mathcal{F}$ be a filter on $\mathbb{N}$ (containing all co-finite sets). Define a topology $\tau_{\mathcal{F}}$ over ${\mathbb{N}}^{<\omega}$ by letting a subset $U$ of ${\mathbb{N}}^{<\omega}$ be open, if $\{n\in\mathbb{N}:\;s\widehat{\;}n\in U\}\in\mathcal{F}$ for all $s\in U$. Then $({\mathbb{N}}^{<\omega},\tau_{\mathcal{F}})$ is Hausdorff, zero-dimensional and without isolated points. Moreover, when $\mathcal{F}$ is a non-principal ultrafilter, $({\mathbb{N}}^{<\omega},\tau_{\mathcal{F}})$ is extremally disconnected, i.e., the closure of an open set is open (see [20] and [6]). Since this space is zero-dimensional and has no isolated points, for every discrete $D\subseteq{\mathbb{N}}^{<\omega}$ there is $s\in{\mathbb{N}}^{<\omega}\setminus D$ such that $D\cup\{s\}$ is still discrete, i.e., there are no maximal discrete subsets of ${\mathbb{N}}^{<\omega}$. It is a classical fact that every countable extremally disconnected Hausdorff space can be embedded in $\beta\mathbb{N}$ (see, for instance, [27, Theorem 1.4.7]). Therefore, if $\mathfrak{E}\subseteq\beta\mathbb{N}$ is homeomorphic to our space, then $\mathfrak{E}$ has no maximal discrete subsets, that is =$\mathcal{I}=\bigcap\{\mathcal{U}^{*}:\mathcal{U}\in\mathfrak{E}\}$ is an $\omega$-maximal ideal which is not strongly $\omega$-maximal (by Theorem 3.4).
Our next result implies that the $\omega$-maximal ideal constructed in the previous example is nowhere maximal.
Proposition 3.9.
Let $\mathfrak{D}=\{\mathcal{I}_{n}\,\colon\,n\in\mathbb{N}\}$ be a collection of maximal ideals and $\mathcal{I}=\bigcap_{n\in\mathbb{N}}\mathcal{I}_{n}$. Then, $\mathfrak{D}$ has an isolated point if and only if there is $A\in\mathcal{I}^{+}$ such that $\mathcal{I}\restriction A$ is a maximal ideal on $A$.
Proof.
Suppose that $\mathcal{I}_{n}^{*}$ is an isolated point of $\mathfrak{D}^{*}$, that is, there is $A\subseteq\mathbb{N}$ such that $\mathfrak{D}^{*}\cap A^{*}=\{\mathcal{I}^{*}_{n}\}$. Then, $A\in\mathcal{I}^{+}$ and $\mathcal{I}\restriction A=\mathcal{I}_{n}\restriction A$ is a maximal ideal on $A$. Conversely, suppose that there is $A\in\mathcal{I}^{+}$ such that $\mathcal{I}\restriction A$ is a maximal ideal on $A$. Notice that $\mathcal{I}\restriction A=\mathcal{I}_{m}\restriction A$ for some unique $m\in\mathbb{N}$. Thus, $\mathcal{I}_{m}^{*}$ is an isolated point of $\mathfrak{D}^{*}$. ∎
4. Dimension of $c_{0,\mathcal{J}}/c_{0,\mathcal{I}}$
In this section we present several results about the dimension of the quotient $c_{0,\mathcal{J}}/c_{0,\mathcal{I}}$ when $\mathcal{I}$ and $\mathcal{J}$ are ideals on $\mathbb{N}$ with $\mathcal{I}\subsetneq\mathcal{J}$. In particular, we obtain some information about $\ell_{\infty}/c_{0,\mathcal{I}}$ which corresponds to the case $\mathcal{J}=\mathcal{P}(\mathbb{N})$.
Now we introduce a relativized version of $k$-maximality.
Definition 4.1.
Let $\mathcal{I}$ and $\mathcal{J}$ be ideals on $\mathbb{N}$ with $\mathcal{I}\subsetneq\mathcal{J}$ and $k$ be a positive integer. We say that $\mathcal{I}$ is $k$-maximal in $\mathcal{J}$ if there is a family $\{\mathcal{L}_{1},\ldots,\mathcal{L}_{k}\}$ of pairwise distinct maximal ideals on $\mathbb{N}$ with $\mathcal{J}\not\subseteq\mathcal{L}_{i}$ for all $1\leq i\leq k$ and $\mathcal{I}=(\bigcap_{i=1}^{k}\mathcal{L}_{i})\cap\mathcal{J}$.
It is clear that, for a positive integer $k$, an ideal is $k$-maximal in the trivial ideal $\mathcal{P}(\mathbb{N})$ if and only if it is $k$-maximal as it was defined in Section 3. To provide a characterization of the relativized version of $k$-maximality we need to introduce a special type of disjoint families.
Definition 4.2.
Let $\mathcal{I}$ and $\mathcal{J}$ be ideals on $\mathbb{N}$ with $\mathcal{I}\subsetneq\mathcal{J}$ and $k$ be a positive integer. A collection $\{A_{1},\ldots,A_{k}\}$ of subsets of $\mathbb{N}$ is called a $(\mathcal{I},\mathcal{J},k)$-family if
  1. (1)
    $\{A_{1},\ldots,A_{k}\}\subseteq\mathcal{J}\setminus\mathcal{I}$;
  2. (2)
    $A_{i}\cap A_{j}=\emptyset$ if $i\neq j$;
  3. (3)
    $B\setminus(A_{1}\cup\cdots\cup A_{k})\in\mathcal{I}$ for all $B\in\mathcal{J}\setminus\mathcal{I}$;
  4. (4)
    $\mathcal{I}\restriction A_{i}$ is a maximal ideal on $A_{i}$ for each $i\in\{1,\ldots,k\}$.
Proposition 4.3.
Let $\mathcal{I}$ and $\mathcal{J}$ be ideals on $\mathbb{N}$ with $\mathcal{I}\subsetneq\mathcal{J}$ and $k$ a positive integer. Then $\mathcal{I}$ is $k$-maximal in $\mathcal{J}$ if and only if there exists a $(\mathcal{I},\mathcal{J},k)$-family.
Proof.
Suppose there is a $(\mathcal{I},\mathcal{J},k)$-family $\{A_{1},\ldots,A_{k}\}$. Define $\mathcal{L}_{i}=(\mathcal{I}\restriction A_{i})\sqcup\mathcal{P}(A_{i}^{c})$ for each $i=1,\ldots,k$. By Lemma 2.2 each $\mathcal{L}_{i}$ is maximal on $\mathbb{N}$ and $A_{i}\in\mathcal{J}\setminus\mathcal{L}_{i}$ for all $1\leq i\leq k$. Since $A_{i}\in\mathcal{L}_{j}$ for all $j\neq i$, we have $\mathcal{L}_{i}\neq\mathcal{L}_{j}$ for $i\neq j$. We claim that $\mathcal{I}=(\bigcap_{i=1}^{k}\mathcal{L}_{i})\cap\mathcal{J}$. The inclusion $\subseteq$ is clear. Now, suppose $B\in(\bigcap_{i=1}^{k}\mathcal{L}_{i})\cap\mathcal{J}$. Then $\bigcup_{i=1}^{k}(B\cap A_{i})\in\mathcal{I}$. As $B=B\setminus(A_{1}\cup\cdots\cup A_{k})\cup(B\cap(A_{1}\cup\cdots\cup A_{k}))$. By condition (3), $B\in\mathcal{I}$.
Conversely, suppose $\mathcal{I}$ is $k$-maximal in $\mathcal{J}$ and let $\{\mathcal{L}_{1},\ldots,\mathcal{L}_{k}\}$ be a family of maximal ideals on $\mathbb{N}$ with $\mathcal{J}\not\subset\mathcal{L}_{i}$ for all $1\leq i\leq k$ and $\mathcal{I}=(\bigcap_{i=1}^{k}\mathcal{L}_{i})\cap\mathcal{J}$. By Theorem 3.1, there is a partition $\{D_{1},\ldots,D_{k}\}$ of $\mathbb{N}$ such that $D_{i}\in\mathcal{L}^{*}_{i}\cap\mathcal{L}_{j}$ for all $j,i$ in $\{1,\ldots,k\}$ with $j\neq i$. For each $1\leq i\leq k$, let $X_{i}\in\mathcal{J}\setminus\mathcal{L}_{i}$. Set $A_{i}=X_{i}\cap D_{i}$ for $i\in\{1,\ldots,k\}$. Then $A_{i}\in\mathcal{J}\setminus\mathcal{L}_{i}$ for all $i$. We will prove that $\{A_{1},\ldots,A_{k}\}$ is a $(\mathcal{I},\mathcal{J},k)$-family. Clearly, the conditions (1) and (2) of Definition 4.2 hold. It remains to check conditions (3) and (4):
  1. (3)
    Let $B\in\mathcal{J}\setminus\mathcal{I}$. Observe that $B\setminus A_{i}\in\mathcal{L}_{i}$ for all $i\in\{1,\ldots,k\}$. So, $B\setminus(A_{1}\cup\ldots\cup A_{k})\in\mathcal{I}$.
  2. (4)
    Fix $i\in\{1,\ldots,k\}$. Then $\mathcal{I}\restriction A_{i}=\mathcal{L}_{i}\restriction A_{i}$. Hence, $\mathcal{I}\restriction A_{i}$ is maximal on $A_{i}$.∎
Remark 4.4.
$(\mathcal{I},\mathcal{J},k)$-families are unique in the following sense. Let $\mathcal{I}$ and $\mathcal{J}$ be ideals on $\mathbb{N}$ with $\mathcal{I}\subsetneq\mathcal{J}$. Suppose that $\mathcal{I}$ is $k$-maximal in $\mathcal{J}$ and $\{A_{1},\ldots,A_{k}\}$ is a $(\mathcal{I},\mathcal{J},k)$-family. Then
  1. (i)
    $\mathcal{J}=\mathcal{I}\sqcup\mathcal{P}(A_{1}\cup\ldots\cup A_{k})$.
  2. (ii)
    If $\{A_{1}^{\prime},\ldots,A_{k}^{\prime}\}$ is also a $(\mathcal{I},\mathcal{J},k)$-family, then $(\bigcup_{1\leq i\leq k}A_{i})\triangle(\bigcup_{1\leq i\leq k}A_{i}^{\prime})\in\mathcal{I}$.
Lemma 4.5.
Let $\mathcal{I}$ and $\mathcal{J}$ be ideals on $\mathbb{N}$ with $\mathcal{I}\subseteq\mathcal{J}$. If $\mathcal{A}\subseteq\mathcal{J}$ is a $\mathcal{I}$-AD family, then the set $\{\chi_{A}+c_{0,\mathcal{I}}\,\colon\,A\in\mathcal{A}\}$ is a linearly independent subset of $c_{0,\mathcal{J}}/c_{0,\mathcal{I}}$.
Proof.
[…]
Theorem 4.6.
Let $\mathcal{I}$ and $\mathcal{J}$ be ideals on $\mathbb{N}$ with $\mathcal{I}\subsetneq\mathcal{J}$ and $k$ be a positive integer. Then $\dim(c_{0,\mathcal{J}}/c_{0,\mathcal{I}})=k$ if and only if $\mathcal{I}$ is $k$-maximal in $\mathcal{J}$.
Proof.
[…]
  1. (1)
    Clearly, $A_{i}\in\mathcal{J}\setminus\mathcal{I}$ for all $i\in\{1,\ldots,k\}$.
  2. (2)
    Since $A_{i}\cap A_{j}\in\mathcal{I}$ for all $i,j\in\{1,\ldots,k\}$ and $i\neq j$, by a standard procedure we can make them pairwise disjoint and thus we may assume that $A_{i}\cap A_{j}=\emptyset$ for each $i,j\in\{1,\ldots,k\}$ with $i\neq j$.
  3. (3)
    Let $B\in\mathcal{J}\setminus\mathcal{I}$ be given and $D\coloneqq B\setminus(A_{1}\cup\cdots\cup A_{k})$. Then there are $\alpha_{1},\ldots,\alpha_{k}\in\mathbb{R}$ such that ${\bf z}=\chi_{D}-\sum_{1\leq j\leq k}\alpha_{j}\chi_{A_{j}}\in c_{0,\mathcal{I}}$. Whence, $D\subseteq A(1/2,{\bf z})\in\mathcal{I}$.
  4. (4)
    Fix $1\leq i\leq k$ and let $B\subseteq A_{i}$. Since $B\in\mathcal{J}$, there exist $\beta_{1},\ldots,\beta_{k}\in\mathbb{R}$ such that ${\bf w}=\chi_{B}-\sum_{1\leq j\leq k}\beta_{j}\chi_{A_{j}}\in c_{0,\mathcal{I}}$. If $\beta_{i}=0$, then $B\subseteq A(1/2,{\bf w})\in\mathcal{I}$. If $\beta_{i}\neq 0$, then $A_{i}\setminus B\subseteq A(|\beta_{i}|,{\bf w})\in\mathcal{I}$. Thus, $\mathcal{I}\restriction A_{i}$ is maximal on $A_{i}$.
Conversely, suppose that $\mathcal{I}$ is $k$-maximal in $\mathcal{J}$ and let $\mathcal{L}_{1},\ldots,\mathcal{L}_{k}$ be maximal ideals on $\mathbb{N}$ such that $\mathcal{I}=(\bigcap_{1\leq i\leq k}\mathcal{L}_{j})\cap\mathcal{J}$. Define $\Phi\colon c_{0,\mathcal{J}}\to\mathbb{R}^{k}$ by $\Phi({\bf x})=(\mathcal{L}_{1}^{*}-\lim{\bf x},\ldots,\mathcal{L}_{k}^{*}-\lim{\bf x})$ if ${\bf x}\in c_{0,\mathcal{J}}$. Notice that $\Phi$ is linear and continuous.
We claim that $\ker\Phi=c_{0,\mathcal{I}}$. If $\Phi({\bf x})={\bf 0}$, then $\mathcal{L}_{j}^{*}-\lim{\bf x}=0$ for all $1\leq j\leq k$. Thus, $A(\varepsilon,{\bf x})\in(\bigcap_{1\leq i\leq k}\mathcal{L}_{j})\cap\mathcal{J}=\mathcal{I}$ for each $\varepsilon>0$. Whence, $\ker\Phi\subset c_{0,\mathcal{I}}$. The converse is clear.
Finally, we show that $\Phi$ is onto. Indeed, there are $A_{1},\ldots,A_{k}\in\mathcal{J}$ satisfying $A_{i}\in\mathcal{L}_{i}^{+}\cap(\bigcap_{1\leq j\neq i\leq k}\mathcal{L}_{j})$ for all $i\in\{1,\ldots,k\}$. So, $\Phi(\chi_{A_{i}})=e_{i}$ for each $1\leq i\leq k$. Therefore, $\Phi$ is onto and we conclude that $\dim c_{0,\mathcal{J}}/c_{0,\mathcal{I}}=k$. ∎
From Theorem 4.6 we obtain the following result.
Corollary 4.7.
Let $\mathcal{I}$ be a proper ideal on a set $\mathbb{N}$ and $k$ be a positive integer. Then $\dim(\ell_{\infty}/c_{0,\mathcal{I}})=k$ if and only if $\mathcal{I}$ is the intersection of exactly $k$ maximal ideals on $\mathbb{N}$.
5. When is $c_{0,\mathcal{I}}$ complemented in $\ell_{\infty}$?
In this section we address the problem of when an ideal $\mathcal{I}$ is complemented, that is, when $c_{0,\mathcal{I}}$ is complemented in $\ell_{\infty}$. We recall that a classical theorem of Lindenstrauss [19], which states that a closed subspace $X\subseteq\ell_{\infty}$ is complemented if and only if $X$ is isomorphic to $\ell_{\infty}$. As it was noticed by Kania in [14], if $\mathcal{I}$ is the intersection of finitely many maximal ideals, then $\mathcal{I}$ is complemented (for a proof see [23, Proposisition 5.20]). Therefore, if $\mathcal{I}$ is the intersection of finitely many maximal ideals on an infinite set $A$, then $c_{0,\mathcal{I}}$ is isomorphic to $\ell_{\infty}(A)$ (and hence, isomorphic to $\ell_{\infty}$).
Recall that a $\mathcal{A}\subseteq\mathcal{I}^{+}$ is $\mathcal{I}$-AD if $X\cap Y\in\mathcal{I}$ for all $X,Y\in\mathcal{A}$. We let
[…]
We will write just $\mathfrak{ad}(\mathcal{A})$ when it is clear from the context which ideal $\mathcal{I}$ is used.
We begin by stating a lemma which is a well-known consequence of Theorem 2.1 (for a proof, see, for instance, [18]).
Lemma 5.1.
Let $\mathcal{I}$ be a meager ideal on $\mathbb{N}$. Then $\mathfrak{ad}(\mathcal{\mathcal{I}}^{+})=2^{\aleph_{0}}$.
On the other hand, $\mathfrak{ad}(\mathcal{\mathcal{I}}^{+})\leq\aleph_{0}$ for every $\omega$-maximal ideal $\mathcal{I}$, this was implicitly shown by Plewik [22] and, for the sake of completion, we include a direct proof. Let $\mathcal{I}=\bigcap_{n}\mathcal{I}_{n}$ be an $\omega$-maximal ideal and suppose $\mathcal{B}\subseteq\mathcal{I}^{+}$ is an uncountable $\mathcal{I}$-AD family. Then, there is $n$ such that $|\mathcal{B}\cap\mathcal{I}_{n}^{+}|\geq 2$. If $A,B\in\mathcal{B}\cap\mathcal{I}_{n}^{+}$ are two different sets, then $A\cap B\in\mathcal{I}\subseteq\mathcal{I}_{n}$, which contradicts that $\mathcal{I}_{n}$ is maximal. Additionally, any $\omega$-maximal ideal is not measurable with respect to the usual product measure on $2^{\mathbb{N}}$ (see [15, section 4.1]).
[…]
Our first result about complemented ideals is based on Lindenstrauss’s theorem and a result from [23].
Theorem 5.2.
Let $\{K_{n}\,\colon\,n\in\mathbb{N}\}$ be a partition of $\mathbb{N}$ and let $\mathcal{I}_{n}$ be an ideal on $K_{n}$ for each $n\in\mathbb{N}$. Suppose that $c_{0,\mathcal{I}_{n}}$ is complemented in $\ell_{\infty}(K_{n})$ for each $n\in\mathbb{N}$. Then, the ideal $\mathcal{J}=\bigoplus_{n\in\mathbb{N}}\mathcal{I}_{n}$ is complemented. In particular, if $\mathcal{I}$ is a strongly $\omega$-maximal ideal, then $\mathcal{I}$ is complemented.
Proof.
By the aforementioned Lindenstrauss’s theorem, $c_{0,\mathcal{I}_{m}}$ is isomorphic to $\ell_{\infty}(K_{m})$ for each $m\in\mathbb{N}$. From [23, Theorem 5.2] it follows that
$c_{0,\mathcal{J}}\cong\ell_{\infty}((c_{0,\mathcal{I}_{m}})_{m\in\mathbb{N}})\sim\ell_{\infty}((\ell_{\infty}(K_{m}))_{m\in\mathbb{N}})\sim\ell_{\infty}.$
Again, by Lindenstrauss’s theorem, we conclude that $c_{0,\mathcal{J}}$ is complemented in $\ell_{\infty}$.
Suppose $\mathcal{I}$ is strongly $\omega$-maximal. By Proposition 3.1, there is a partition $\{K_{n}\}$ of $\mathbb{N}$ such that $\mathcal{I}=\bigoplus_{n\in\mathbb{N}}(\mathcal{I}_{n}\restriction K_{n})$ and $\mathcal{I}_{n}\restriction K_{n}$ is maximal on $K_{n}$ for each $n$. Then, the previous argument shows that $c_{0,\mathcal{I}}$ is isomorphic to $\ell_{\infty}$. ∎
Corollary 5.3.
Let $\mathcal{I}$ be an ideal on $\mathbb{N}$. Then $\mathcal{I}$ is complemented if and only if $\mathcal{I}^{\omega}$ is complemented.
Proof.
It follows from Theorem 5.2 that if $\mathcal{I}$ is complemented, then $\mathcal{I}^{\omega}$ is complemented. Conversely, suppose that $\mathcal{I}^{\omega}$ is complemented. By [23, Theorem 5.2] we have $c_{0,\mathcal{I}^{\omega}}\cong\ell_{\infty}(c_{0,\mathcal{I}})$. Consequently, $c_{0,\mathcal{I}}$ is isomorphic to a complemented subspace of $c_{0,\mathcal{I}^{\omega}}$. On the other hand, notice that $c_{0,\mathcal{I}^{\omega}}\sim\ell_{\infty}$ by Lindenstrauss’s theorem. Thus, $c_{0,\mathcal{I}}$ is isomorphic to a complemented subspace of $\ell_{\infty}$. Therefore, $\mathcal{I}$ is complemented. ∎
Proposition 5.4.
Let $\mathcal{I}$ and $\mathcal{J}$ be ideals on $\mathbb{N}$ with $\mathcal{I}\subsetneq\mathcal{J}$ and $k$ be a positive integer. If $\mathcal{J}$ is complemented and $\mathcal{I}$ is $k$-maximal in $\mathcal{J}$, then $\mathcal{I}$ is complemented.
Proof.
Since $\mathcal{I}$ is $k$-maximal in $\mathcal{J}$, it follows that $c_{0,\mathcal{I}}$ is complemented in $c_{0,\mathcal{J}}$ by Theorem 4.6. Additionally, $c_{0,\mathcal{J}}$ is complemented in $\ell_{\infty}$. Therefore, $c_{0,\mathcal{I}}$ is complemented in $\ell_{\infty}$. This completes the proof. ∎
Remark 5.5.
Despite these results, it remains unknown whether the intersection of a countable family of complemented ideals is itself complemented. Even in the case of two complemented ideals, $\mathcal{I}$ and $\mathcal{J}$, we do not know whether $\mathcal{I}\cap\mathcal{J}$ is complemented.
We say that a proper ideal $\mathcal{I}$ on a set $A\subseteq\mathbb{N}$ is complemented in $A$ if $c_{0,\mathcal{I}}$ is complemented in $\ell_{\infty}(A)$ (recall that $c_{0,\mathcal{I}}$ consists of sequences of the form $(x_{n})_{n\in A}$). The next result shows that the complementation of ideals is preserved by taking restrictions.
Proposition 5.6.
Let $\mathcal{I}$ be an ideal on $\mathbb{N}$ and let $A$ and $B$ be infinite subsets of $\mathbb{N}$ with $A\subseteq B$ and $A\in\mathcal{I}^{+}$. If $\mathcal{I}\restriction B$ is complemented in $B$, then $\mathcal{I}\restriction A$ is complemented in $A$. In particular, if $\mathcal{I}$ is complemented, then $\mathcal{I}\restriction A$ is complemented in $A$ for every $A\in\mathcal{I}^{+}$.
Proof.
Since $c_{0,\mathcal{I}\restriction B}$ is complemented in $\ell_{\infty}(B)$, it suffices to prove that $c_{0,\mathcal{I}\restriction A}$ is isometric to a complemented subspace of $c_{0,\mathcal{I}\restriction B}$. To this end, consider the map $\varphi\colon c_{0,\mathcal{I}\restriction A}\to c_{0,\mathcal{I}\restriction B}$ defined by
$(x_{n})\mapsto(\tilde{x}_{n}),\quad\mbox{where}\quad\tilde{x}_{n}=\begin{cases}x_{n},&n\in A;\\ 0,&n\in B\setminus A.\end{cases}$
This mapping is an isometry. Furthermore, the map $P\colon c_{0,\mathcal{I}\restriction B}\to\varphi(c_{0,\mathcal{I}\restriction A})$ given by
$(x_{n})\mapsto(y_{n}),\quad\mbox{where}\quad y_{n}=\begin{cases}x_{n},&n\in A;\\ 0,&n\in B\setminus A,\end{cases}$
defines a projection, completing the proof. ∎
Now we present a more informative proof of Theorem 5.2. But first we need an auxiliary result. Recall that if ${\bf x}=(x_{n})\in\ell_{\infty}$, its support is defined as $\mathrm{supp}({\bf x})\coloneqq\{n\in\mathbb{N}\,\colon\,x_{n}\neq 0\}$. Let $F\subseteq\mathbb{N}$, and suppose that $({\bf x}_{n})_{n\in F}$ is a sequence in $\ell_{\infty}$ such that $\mathrm{supp}({\bf x}_{n})\cap\mathrm{supp}({\bf x}_{n})=\emptyset$ for all $m\neq n$. Then, the sum $\sum_{n\in F}{\bf x}_{n}$ is the sequence in $\ell_{\infty}$ whose $m$-th coordinate is ${\bf x}_{n}(m)$ whenever $m\in\mathrm{supp}({\bf x}_{n})$ for some $n\in F$, and 0 otherwise.
Lemma 5.7.
Let $\{\mathcal{I}_{n}\,\colon\,n\in\mathbb{N}\}$ be a collection of maximal ideals and let $\{A_{n}\,\colon\,n\in\mathbb{N}\}$ be a collection of pairwise disjoint subsets of $\mathbb{N}$ such that $A_{n}\in\mathcal{I}_{n}^{*}$ for each $n\in\mathbb{N}$. Let $\mathcal{I}=\bigcap_{n\in\mathbb{N}}\mathcal{I}_{n}$. Then
$P\colon\ell_{\infty}$ $\to c_{0,\mathcal{I}}$
${\bf x}$ $\mapsto{\bf x}-\sum_{m\in\mathbb{N}}(\mathcal{I}_{m}^{*}-\lim{\bf x})\chi_{A_{m}}$
is a continuous projection from $\ell_{\infty}$ onto $c_{0,\mathcal{I}}$.
Proof.
Firstly, we prove that $P$ is well defined. Let ${\bf x}=(x_{n})\in\ell_{\infty}$ and ${\bf y}=P({\bf x})=(y_{n})$. Let $\varepsilon>0$ and $m\in\mathbb{N}$, we have
$A(\varepsilon,{\bf y})\cap A_{m}=\{n\in A_{m}\,\colon\,|y_{n}|\geq\varepsilon\}=\{n\in A_{m}\,\colon\,|x_{n}-(\mathcal{I}_{m}^{*}-\lim{\bf x})|\geq\varepsilon\}.$
From the definition of $\mathcal{I}_{m}^{*}-\lim{\bf x}$ it follows that $\{n\in\mathbb{N}\,\colon\,|x_{n}-(\mathcal{I}_{m}^{*}-\lim{\bf x})|\geq\varepsilon\}\in\mathcal{I}_{m}$. Thus, $A(\varepsilon,{\bf y})\cap A_{m}\in\mathcal{I}_{m}$ for each $\varepsilon>0$ and $m\in\mathbb{N}$. By Lemma 2.3, $A(\varepsilon,{\bf y})\in\mathcal{I}$ for every $\varepsilon>0$. We conclude that ${\bf y}=P({\bf x})\in c_{0,\mathcal{I}}$. Therefore, $P$ is well defined.
Clearly, $P$ is linear and $\|P\|\leq 2$. To check that $P$ is a projection, let ${\bf x}\in\ell_{\infty}$ and ${\bf y}=P({\bf x})$. We have
$P(P({\bf y}))=P({\bf y})-\sum_{m\in\mathbb{N}}(\mathcal{I}_{m}^{*}-\lim P({\bf x}))\chi_{A_{m}}=P({\bf y}),$
because of $P({\bf x})\in c_{0,\mathcal{I}_{m}}$, that is, $\mathcal{I}_{m}^{*}-\lim P({\bf x})=0$ for each $m\in\mathbb{N}$. The previous argument also shows that $P(\ell_{\infty})=c_{0,\mathcal{I}}$. Hence, $P$ is a projection from $\ell_{\infty}$ onto $c_{0,\mathcal{I}}$. ∎
Theorem 5.8.
Let $\mathcal{I}$ be a strongly $\omega$-maximal ideal. Then, there is a positive projection $Q\colon\ell_{\infty}\to\ell_{\infty}$ such that $\ker Q=c_{0,\mathcal{I}}$. Moreover, if for each $B\subseteq\mathbb{N}$ we let $T(B)=\{n\in\mathbb{N}\,\colon\,Q(\chi_{B})_{n}=1\}$, then the following properties hold:
  1. (1)
    $Q(\chi_{A})=\chi_{T(A)}$ for all $A\subseteq\mathbb{N}$ and $Q(\chi_{\mathbb{N}})=\chi_{\mathbb{N}}$;
  2. (2)
    If $A\subseteq B$, then $T(A)\subseteq T(B)$;
  3. (3)
    $T(T(A))=T(A)$ for all $A\subseteq\mathbb{N}$;
  4. (4)
    $T(A\cap B)=T(A)\cap T(B)$ for all $A,B\subseteq\mathbb{N}$;
  5. (5)
    The family =$\mathcal{B}=\{B\subseteq\mathbb{N}\,\colon\,T(B)=B\}$ is closed under arbitrary intersections.
Proof.
Let $\{\mathcal{I}_{j}\,\colon\,j\in\mathbb{N}\}$ be a discrete collection of pairwise distinct maximal ideals on $\mathbb{N}$ such that $\mathcal{I}=\bigcap_{j\in\mathbb{N}}\mathcal{I}_{j}$. By Proposition 3.1, there is a partition $\{A_{m}\,\colon\,m\in\mathbb{N}\}$ of $\mathbb{N}$ with $A_{m}\in\mathcal{I}_{m}^{*}$ for each $m\in\mathbb{N}$. Consider the projection defined in Lemma 5.7, that is,
$P\colon\ell_{\infty}$ $\to c_{0,\mathcal{I}}$
${\bf x}$ $\mapsto{\bf x}-\sum_{m\in\mathbb{N}}(\mathcal{I}_{m}^{*}-\lim{\bf x})\chi_{A_{m}}.$
Now define $Q\colon\ell_{\infty}\to\ell_{\infty}$ by $Q=Id-P$, that is,
$Q({\bf x})=\sum_{m\in\mathbb{N}}(\mathcal{I}_{m}^{*}-\lim{\bf x})\chi_{A_{m}},\quad{\bf x}\in\ell_{\infty}.$ (5.1)
Notice that $Q$ is a positive projection and $\ker Q=c_{0,\mathcal{I}}$. Now we will check properties (1)-(5).
Observe that (1) follows from the definition of $Q$. Properties (2) and (3) follow from positiveness and idempotence of $Q$. For (4), notice that if $n\in\mathbb{N}$, then $Q(\chi_{A})_{n}=1$ if and only if there is $m\in\mathbb{N}$ such that $n\in A_{m}$ and $A\in\mathcal{I}_{m}^{*}$. Let $A,B\subseteq\mathbb{N}$ be given. By (2) it suffices to show that $T(A)\cap T(B)\subseteq T(A\cap B)$. Let $n\in T(A)\cap T(B)$, that is, $Q(\chi_{A})_{n}=1$ and $Q(\chi_{B})_{n}=1$. Thus, there is $m\in\mathbb{N}$ such that $n\in A_{m}$, $A\in\mathcal{I}_{m}^{*}$ and $B\in\mathcal{I}_{m}^{*}$. Therefore, $A\cap B\in\mathcal{I}_{m}^{*}$. So, $n\in T(A\cap B)$.
Now we will prove (5).
Claim 5.9.
Let $B\subseteq\mathbb{N}$ be such that $T(B)=B$. If $A_{m}\cap B\neq\emptyset$, then $A_{m}\subseteq B$.
Let $k\in A_{m}$. If $n\in A_{m}\cap B$, then $n\in T(B)$. Thus, there is $j\in\mathbb{N}$ such that $n\in A_{j}$ and $B\in\mathcal{I}_{j}^{*}$. As $n\in A_{m}$, $j=m$. Hence, $k\in A_{m}$ and $B\in\mathcal{I}_{m}^{*}$, that is, $k\in T(B)=B$.
Now let $\{C_{i}\}_{i\in I}$ be a collection in $\mathcal{B}$. By (2) we have $T(\bigcap_{i\in I}C_{i})\subseteq\bigcap_{i\in I}C_{i}$. Let $n\in\bigcap_{i\in I}C_{i}$. For each $i\in\mathbb{N}$, there is $m_{i}\in\mathbb{N}$ such that $n\in A_{m_{i}}$ and $C_{i}\in\mathcal{I}_{m_{i}}^{*}$. Observe that $(m_{i})_{i\in I}$ is a constant family, say $m_{i}=m$ for each $i\in I$. Since $A_{m}\cap C_{i}\neq\emptyset$ and $T(C_{i})=C_{i}$ for each $i\in I$, by Claim 5.9 we obtain $A_{m}\subseteq C_{i}$ for all $i\in I$. Consequently, $A_{m}\subseteq\bigcap_{i\in I}C_{i}$. It follows that $\bigcap_{i\in I}C_{i}\in\mathcal{I}_{m}^{*}$. So, $n\in T(\bigcap_{i\in I}C_{i})$ and we are done. ∎
Remark 5.10.
Property (4) of Theorem 5.8 can be restated as follows: For any pair $A,B$ of subsets of $\mathbb{N}$ and $n\in\mathbb{N}$,
$Q(\chi_{A})_{n}\neq 0\quad\mbox{and}\quad Q(\chi_{B})_{n}\neq 0\;\Rightarrow\;Q(\chi_{A\cap B})_{n}\neq 0.$ (5.2)
We will show that, under some extra condition, the complementation of $\mathcal{I}$ implies that $\mathcal{I}$ is $\omega$-maximal. For that end, we need a key lemma whose argument is motivated by the proof of the non-complementation of $c_{0}$ in $\ell_{\infty}$ due to Whitley (see [7, Theorem 5.6] or [1, Theorem 2.5.4]). Indeed, we will show a more general fact: $c_{0}$ is not complemented in $c_{0,\mathcal{J}}$ for any ideal $\mathcal{J}\supsetneq{\sf Fin}$. A similar result appears in [18, Corollary 1.5] about the space $c$ of convergent sequences.
Lemma 5.11.
Let $\mathcal{I}\subsetneq\mathcal{J}$ be ideals on $\mathbb{N}$ such that $c_{0,\mathcal{I}}$ is complemented in $c_{0,\mathcal{J}}$. Then, there exists a collection $\{\mathcal{A}_{n,k}\colon\,n,k\in\mathbb{N}\}$ of subsets of $2^{\mathbb{N}}$ satisfying:
  1. (1)
    $\mathcal{A}_{n,k}\subseteq\mathcal{J}\cap\mathcal{I}^{+}$ for each $n,k\in\mathbb{N}$.
  2. (2)
    $\mathcal{A}_{n,k}\subseteq\mathcal{A}_{n,k+1}$ for all $n,k\in\mathbb{N}$.
  3. (3)
    There is $M>0$ such that $\mathfrak{ad}_{\mathcal{I}}(\mathcal{A}_{n,k})\leq kM$ for all $n,k\in\mathbb{N}$.
  4. (4)
    For each $n\in\mathbb{N}$, let $\mathcal{A}_{n}=\bigcup_{k\in\mathbb{N}}\mathcal{A}_{n,k}$. Then $\mathcal{J}\cap\mathcal{I}^{+}=\bigcup_{n\in\mathbb{N}}\mathcal{A}_{n}$.
  5. (5)
    If $A\in\mathcal{A}_{n,k}$ and $A=B\cup C$ with $B\cap C\not\in\mathcal{A}_{n}$, then either $B\in\mathcal{A}_{n,2k}$ or $C\in\mathcal{A}_{n,2k}$. In particular, if $A\in\mathcal{A}_{n}$ and $A=B\cup C$ with $B\cap C\not\in\mathcal{A}_{n}$, then either $B\in\mathcal{A}_{n}$ or $C\in\mathcal{A}_{n}$.
Proof.
Let $P\colon c_{0,\mathcal{J}}\to c_{0,\mathcal{I}}$ be a projection onto $c_{0,\mathcal{I}}$ and $Q=Id-P$. For each $n,k\in\mathbb{N}$, define
$\mathcal{A}_{n,k}=\{A\in\mathcal{J}\,\colon\,|Q(\chi_{A})_{n}|\geq 1/k\}.$
Since =$\mathcal{I}=\{A\in\mathcal{J}\,\colon\,Q(\chi_{A})={\bf 0}\}$, we have
$\mathcal{J}\cap\mathcal{I}^{+}$ =$=\{A\in\mathcal{J}\,\colon\,Q(\chi_{A})\neq{\bf 0}\}$
$=\bigcup_{k,n\in\mathbb{N}}\{A\in\mathcal{J}\,\colon\,|Q(\chi_{A})_{n}|\geq 1/k\}$
$=\bigcup_{n\in\mathbb{N}}\mathcal{A}_{n}.$
Then (1) and (4) hold. (2) is obvious. To see (3), let $m\in\mathbb{N}$ be given and $A_{1},\ldots,A_{m}\in\mathcal{A}_{n,k}$ be such that $A_{i}\cap A_{j}\in\mathcal{I}$ for $i\neq j$. If $F_{1}=A_{1}$ and $F_{j}=A_{j}\setminus(A_{1}\cup\cdots\cup A_{j-1})$ for $2\leq j\leq m$, then $Q(\chi_{A_{j}})=Q(\chi_{F_{j}})$ for all $1\leq j\leq m$. For $1\leq j\leq m$, let $a_{j}\in\mathbb{K}$ be such that $a_{j}Q(\chi_{F_{j}})_{n}=|Q(\chi_{F_{j}})_{n}|$. Thus,
$\frac{m}{k}\leq\sum_{j=1}^{m}|Q(\chi_{A_{j}})_{n}|$ $=\sum_{j=1}^{m}|Q(\chi_{F_{j}})_{n}|$
$=\sum_{j=1}^{m}a_{j}Q(\chi_{F_{j}})_{n}$
$\leq\left\|\sum_{j=1}^{m}a_{j}Q(\chi_{F_{j}})\right\|$
$\leq\|Q\|\left\|\sum_{j=1}^{m}a_{j}\chi_{F_{j}}\right\|=\|Q\|.$
Hence, $m\leq k\|Q\|$.
Finally, we show (5). Let $A\in\mathcal{A}_{n,k}$ and $A=B\cup C$ with $B\cap C\not\in\mathcal{A}_{n}$. Then $|Q(\chi_{B\cap C})_{n}|=0$. Assume that $|Q(\chi_{B})_{n}|<1/2k$ and $|Q(\chi_{C})_{n}|<1/2k$. Note that $Q(\chi_{B})_{n}=Q(\chi_{B\setminus C})_{n}$ and $Q(\chi_{C})_{n}=Q(\chi_{C\setminus B})_{n}$. Thus
$|Q(\chi_{A})_{n}|=|Q(\chi_{A\setminus B\cap C})_{n}|=|Q(\chi_{B})_{n}+Q(\chi_{C})_{n}|\leq 1/k,$
which is absurd. ∎
Theorem 5.12.
Let $\mathcal{I}\subsetneq\mathcal{J}$ be proper ideals on $\mathbb{N}$ such that $c_{0,\mathcal{I}}$ is complemented in $c_{0,\mathcal{J}}$. Then $\mathfrak{ad}_{\mathcal{I}}(\mathcal{J}\cap\mathcal{I}^{+})\leq\aleph_{0}$.
Proof.
For $k,n\in\mathbb{N}$, let $\mathcal{A}_{n,k}$ be as in Lemma 5.11. Then, we have $\mathcal{J}\cap\mathcal{I}^{+}=\bigcup_{n,k}\mathcal{A}_{n,k}$ and $\mathfrak{ad}_{\mathcal{I}}(\mathcal{A}_{n,k})<\aleph_{0}$ for each $n,k$. Suppose $\mathcal{B}\subseteq\mathcal{J}\cap\mathcal{I}^{+}$ is an $\mathcal{I}$-AD family. Then, for every $n,k\in\mathbb{N}$, $\mathcal{B}\cap\mathcal{A}_{n,k}$ is finite. Thus, $\mathcal{B}$ is countable, and consequently $\mathfrak{ad}_{\mathcal{I}}(\mathcal{J}\cap\mathcal{I}^{+})\leq\aleph_{0}$. ∎
The next observation follows immediately from Lemma 5.1.
Corollary 5.13.
Let $\mathcal{I}\subsetneq\mathcal{J}$ be proper ideals on $\mathbb{N}$ such that $\mathcal{I}\restriction A$ is meager as a subset of $2^{A}$ for some infinite set $A\in\mathcal{J}$. Then $\mathfrak{ad}_{\mathcal{I}}(\mathcal{J}\cap\mathcal{I}^{+})>\aleph_{0}$.
An ideal $\mathcal{I}$ is everywhere meager [8], if $\mathcal{I}\restriction A$ is meager in $2^{A}$ for all $A\in\mathcal{I}^{+}$, which in turns is equivalent to requiring that $\mathcal{I}\restriction A$ is Baire measurable in $2^{A}$ for all $A\in\mathcal{I}^{+}$. Every analytic ideal is everywhere meager. On the opposite side, a complemented ideal $\mathcal{I}$ is nowhere meager, since $\mathcal{I}\restriction A$ is not meager in $2^{A}$ for all $A\in\mathcal{I}^{+}$ by Proposition 5.6.
From Theorem 5.12 we get the following.
Corollary 5.14.
Let $\mathcal{I}\subsetneq\mathcal{J}$ be proper ideals on $\mathbb{N}$ such that $\mathcal{I}$ is everywhere meager. Then $c_{0,\mathcal{I}}$ is not complemented in $c_{0,\mathcal{J}}$. In particular, $c_{0}$ is not complemented in $c_{0,\mathcal{J}}$ for any $\mathcal{J}\supsetneq{\sf Fin}$.
Remark 5.15.
Concerning the above results, Theorem 4.6 gives examples of proper ideals $\mathcal{I}$ and $\mathcal{J}$ with ${\sf Fin}\subsetneq\mathcal{I}\subsetneq\mathcal{J}$ such that $c_{0,\mathcal{I}}$ is complemented in $c_{0,\mathcal{J}}$. Indeed, if $\mathcal{J}$ is a proper ideal and $\mathcal{I}_{1},\ldots,\mathcal{I}_{k}$ are maximal ideals such that $\mathcal{J}\not\subseteq\mathcal{I}_{j}$ for all $1\leq j\leq k$, then $\mathcal{I}\coloneqq\mathcal{J}\cap(\bigcap_{1\leq j\leq k}\mathcal{I}_{j})$ is $k$-maximal in $\mathcal{J}$. Thus, $c_{0,\mathcal{I}}$ is complemented in $c_{0,\mathcal{J}}$ by Theorem 4.6. Notice that in this case $\dim(c_{0,\mathcal{J}}/c_{0,\mathcal{I}})=k$. We do not know if there exist examples with $\dim(c_{0,\mathcal{J}}/c_{0,\mathcal{I}})=\infty$.
In spite of the previous results, it is still possible for $c_{0,\mathcal{I}}$ to contain complemented subspaces that are isomorphic or isometric to $c_{0}$. In particular, we proved in [23] that $\ell_{\infty}(c_{0})$ is isometric to $c_{0,{\sf Fin}^{\omega}}$, and since $\ell_{\infty}(c_{0})$ clearly contains a complemented copy of $c_{0}$, the same holds for $c_{0,{\sf Fin}^{\omega}}$. This contrasts with the situation in $\ell_{\infty}$, where every subspace isomorphic or isometric to $c_{0}$ is necessarily not complemented. This leads to the following result.
Proposition 5.16.
Let $\mathcal{I}$ be an ideal on $\mathbb{N}$. If $c_{0,\mathcal{I}}$ has a complemented copy of $c_{0}$, then $\mathcal{I}$ is not complemented.
Proof.
Suppose that $\mathcal{I}$ is complemented, that is, $\ell_{\infty}=c_{0,\mathcal{I}}\oplus W$. Since $c_{0,\mathcal{I}}$ has a complemented copy of $c_{0}$, there is a subspace $E$ isomorphic to $c_{0}$ such that $c_{0,\mathcal{I}}=E\oplus Z$. Thus, $E$ is complemented in $\ell_{\infty}$, which is impossible by Lindenstrauss’s theorem. ∎
However, it is unclear whether $c_{0,\mathcal{I}}$ contains a complemented copy of $c_{0}$ for any meager ideal $\mathcal{I}$.
The next corollary is due to Kania [14, Theorem A].
Corollary 5.17.
Let $\mathcal{I}$ be an ideal on $\mathbb{N}$. If $\mathcal{I}$ is complemented, then $\mathfrak{ad}(\mathcal{I}^{+})\leq\aleph_{0}$.
Now we prove, with an extra assumption, that the complementation of $c_{0,\mathcal{I}}$ implies that $\mathcal{I}$ is $\omega$-maximal.
Theorem 5.18.
Let $\mathcal{I}$ be a proper ideal on $\mathbb{N}$. Suppose $c_{0,\mathcal{I}}$ is complemented in $\ell_{\infty}$ by a projection map $P\colon\ell_{\infty}\to c_{0,\mathcal{I}}$ such that $Q:=Id-P$ satisfies (5.2). Then $\mathcal{I}$ is $\omega$-maximal.
Proof.
Suppose $P\colon\ell_{\infty}\to c_{0,\mathcal{I}}$ is a projection map as in the hypothesis. For each $n\in\mathbb{N}$ consider the family
$\mathcal{A}_{n}=\{A\subseteq\mathbb{N}:\;Q(\chi_{A})_{n}\neq 0\}$
as in Lemma 5.11 with $\mathcal{J}=\mathcal{P}(\mathbb{N})$. Observe that $\mathcal{I}^{+}=\bigcup_{n\in\mathbb{N}}\mathcal{A}_{n}$ by Lemma 5.11. For each $n\in\mathbb{N}$, let
$\mathcal{H}_{n}=\{B\subseteq\mathbb{N}:(\exists\,A\in\mathcal{A}_{n})(A\subseteq B)\}.$
Notice that $\mathcal{H}_{n}\subseteq\mathcal{I}^{+}$ for all $n$. Indeed, suppose not and let $B\in\mathcal{H}_{n}$ with $B\not\in\mathcal{I}^{+}$. Let $A\in\mathcal{A}_{n}$ be such that $A\subseteq B$. Since $B\in\mathcal{I}$, $A\in\mathcal{I}$ and thus $Q(\chi_{A})={\bf 0}$, which is absurd.
We claim that $\mathcal{I}_{n}\coloneqq 2^{\mathbb{N}}\setminus\mathcal{H}_{n}$ is an ideal. By construction, $\mathcal{I}_{n}$ is closed by taking subsets. Now, let $B,C\in\mathcal{I}_{n}$ and suppose that $B\cup C\not\in\mathcal{I}_{n}$, that is, $B\cup C\in\mathcal{H}_{n}$. Let $A\in\mathcal{A}_{n}$ be such that $A\subseteq B\cup C.$ Since $B\cap C\not\in\mathcal{A}_{n}$, by (5) of Lemma 5.11, we have either $A\cap B\in\mathcal{A}_{n}$ or $A\cap C\in\mathcal{A}_{n}$. It follows that either $B\in\mathcal{H}_{n}$ or $C\in\mathcal{H}_{n}$, a contradiction. So, $\mathcal{I}_{n}$ is an ideal. Now as $\mathcal{I}$ is proper, $\mathbb{N}\not\in\mathcal{I}$. Thus, $S=\{n\in\mathbb{N}\,\colon\,Q(\chi_{\mathbb{N}})_{n}\neq 0\}$ is non-empty. Observe that $\mathbb{N}\in\mathcal{A}_{n}$ if and only if $n\in S$. Therefore, $\mathcal{I}_{n}\neq\mathcal{P}(\mathbb{N})$ if and only if $n\in S$.
By (4) of Lemma 5.11, $\mathcal{I}=\bigcap_{n}\mathcal{I}_{n}=\bigcap_{n\in S}\mathcal{I}_{n}$. It remains to show that each $\mathcal{I}_{n}$ is maximal for each $n\in S$. If not, let $B\subseteq\mathbb{N}$ be such that $B\not\in\mathcal{I}_{n}$ and $\mathbb{N}\setminus B\not\in\mathcal{I}_{n}$. Then there are $C_{0},C_{1}\in\mathcal{A}_{n}$ such that $C_{0}\subseteq B$ and $C_{1}\subseteq\mathbb{N}\setminus B$. So, $Q(\chi_{C_{0}\cap C_{1}})={\bf 0}$. On the other hand, by (5.2) we have $Q(\chi_{C_{0}\cap C_{1}})_{n}\neq 0$, an absurd. ∎
Remark 5.19.
It is not difficult to verify that $\mathcal{I}_{n}$ is maximal if and only if Property (5.2) holds for that $n$. We do not know whether the assumption of Property (5.2) can be removed.
Theorem 5.18 raises the following question: Is there an $\omega$-maximal ideal that is not complemented? In other words, is being complemented equivalent to being $\omega$-maximal? A natural candidate for a negative answer to this question is the $\omega$-maximal ideal constructed in Example 3.8.
Our next result is the converse of Theorem 5.8.
Theorem 5.20.
Let $\mathcal{I}$ be a proper ideal on $\mathbb{N}$. Assume that there is an operator $Q\colon\ell_{\infty}\to\ell_{\infty}$ such that
  • $Q$ is a positive projection such that $\ker Q=c_{0,\mathcal{I}}$.
  • For each $A\subseteq\mathbb{N}$, there exists $T(A)\subseteq\mathbb{N}$ such that $Q(\chi_{A})=\chi_{T(A)}$.
  • If $A,B\subseteq\mathbb{N}$, then $T(A\cap B)=T(A)\cap T(B)$.
  • The family =$\mathcal{B}=\{B\subseteq\mathbb{N}\,\colon\,T(B)=B\}$ is closed under arbitrary intersections.
Then $\mathcal{I}$ is strongly $\omega$-maximal. Moreover, there is a partition of $\mathbb{N}$ such that $Q$ has the same form as given in (5.1).
Proof.
For each $n\in\mathbb{N}$, let $V_{n}=\bigcap\{B\in\mathcal{B}\,\colon\,n\in B\}$. Since $\mathcal{B}$ is closed under arbitrary intersections, it follows that $V_{n}\in\mathcal{B}$, i.e., $T(V_{n})=V_{n}$ for every $n\in\mathbb{N}$.
Claim 5.21.
The map $T\colon A\in\mathcal{P}(\mathbb{N})\mapsto T(A)\in\mathcal{P}(\mathbb{N})$ has the following properties:
  1. (1)
    For each $A\subseteq\mathbb{N}$, $T(T(A))=T(A)$;
  2. (2)
    If $A\subseteq B$, then $T(A)\subseteq T(B)$.
  3. (3)
    $Q(\chi_{\mathbb{N}})=\chi_{\mathbb{N}}$.
Proof of the Claim 5.21.
Properties (1) and (2) follow from idempotence and positiveness of $Q$. For (3), it suffices to show that $\mathbb{N}\subseteq T(\mathbb{N})$. Let $n\in\mathbb{N}$ be given. As $n\in V_{n}$ and $V_{n}=T(V_{n})\subseteq T(\mathbb{N})$, we have $n\in T(\mathbb{N})$. ∎
We will follow the notation used in the proof of Theorem 5.18. Recall that for each $n\in\mathbb{N}$, we set $\mathcal{A}_{n}=\{A\subseteq\mathbb{N}\,\colon Q(\chi_{A})_{n}\neq 0\}$. As $Q$ is a positive operator, $\mathcal{H}_{n}=\mathcal{A}_{n}$ for all $n\in\mathbb{N}$. Also notice that if $A,B\subseteq\mathbb{N}$ satisfy that $Q(\chi_{A})_{n}=\chi_{T(A)}(n)\neq 0$ and $Q(\chi_{B})_{n}=\chi_{T(B)}(n)\neq 0$, then $n\in T(A)\cap T(B)=T(A\cap B)$, that is, $Q(\chi_{A\cap B})_{n}=1$. Thus, $Q$ verifies the property (5.2). Hence, by the proof of Theorem 5.18, we have $\mathcal{I}_{n}=2^{\mathbb{N}}\setminus\mathcal{A}_{n}$ is a maximal ideal for each $n\in\{m\in\mathbb{N}\,\colon\,Q(\chi_{\mathbb{N}})_{m}\neq 0\}=\mathbb{N}$ (using part (3) of Claim 5.21). Also, we have that $\mathcal{I}=\bigcap_{n\in\mathbb{N}}\mathcal{I}_{n}$.
Claim 5.22.
For all $n,m\in\mathbb{N}$, $\mathcal{A}_{m}\subseteq\mathcal{A}_{n}$ if and only if $V_{n}\subseteq V_{m}$. In particular, for each $m,n\in\mathbb{N}$ it holds $\mathcal{A}_{m}=\mathcal{A}_{n}$ if and only if $V_{n}=V_{m}$.
Proof of the Claim 5.22.
Suppose $\mathcal{A}_{m}\subseteq\mathcal{A}_{n}$. Since $V_{m}\in\mathcal{A}_{m}$, we have $V_{m}\in\mathcal{A}_{n}$, that is, $n\in T(V_{m})=V_{m}$. Thus, $V_{n}\subseteq V_{m}$.
Conversely, suppose $V_{n}\subseteq V_{m}$. Let $B\in\mathcal{A}_{m}$, that is, $m\in T(B)$. Since $T(T(B))=T(B)$, $T(B)\in\mathcal{B}$. Thus, $V_{m}\subseteq T(B)$. Hence, $V_{n}\subseteq T(B)$. In particular, $n\in T(B)$, that is, $B\in\mathcal{A}_{n}$. Therefore, $\mathcal{A}_{n}\subseteq\mathcal{A}_{m}$. ∎
We will show that if $V_{n}\cap V_{m}\neq\emptyset$, then $V_{n}=V_{m}$, for all $m,n\in\mathbb{N}$. Indeed, let $k\in V_{n}\cap V_{m}$. As $V_{n}\cap V_{m}\in\mathcal{B}$, we have that $V_{k}\subseteq V_{m}\cap V_{n}$. By Claim 5.22, $\mathcal{A}_{n}\subseteq\mathcal{A}_{k}$. Thus by the maximality of $\mathcal{A}_{n}$, $\mathcal{A}_{n}=\mathcal{A}_{k}$. Analogously, $\mathcal{A}_{m}=\mathcal{A}_{k}$. Hence $V_{n}=V_{m}$.
Now consider the following equivalence relation on $\mathbb{N}$: $m\sim n$ if and only if $\mathcal{A}_{n}=\mathcal{A}_{m}$. By Claim 5.22 we have $n\sim m$ if and only if $V_{n}=V_{m}$. Moreover, we have shown that $n\not\sim m$ if and only if $V_{n}\cap V_{m}=\emptyset$. Let $F$ be a complete set of representatives and consider the collection $\{V_{n}\,\colon\,n\in F\}$. Notice that if $n,m\in F$ and $n\neq m$, then $V_{n}\cap V_{m}=\emptyset$. Also, $V_{n}\in\mathcal{I}_{n}^{*}=\mathcal{A}_{n}$ for each $n\in F$. Thus $\{\mathcal{I}_{n}\,\colon\,n\in F\}$ is discrete. Moreover, since $\mathcal{I}=\bigcap_{n\in\mathbb{N}}\mathcal{I}_{n}$ and $F$ is a complete set of representatives, we have $\mathcal{I}=\bigcap_{n\in F}\mathcal{I}_{n}$. Consequently, $\mathcal{I}$ is strongly $\omega$-maximal.
Finally, we will prove that
$Q({\bf x})=\sum_{m\in F}(\mathcal{I}_{m}^{*}-\lim x)\chi_{V_{m}},\quad\text{for all ${\bf x}\in\ell_{\infty}$.}$
For all $A\subseteq\mathbb{N}$ and $n\in\mathbb{N}$ it holds that $\mathcal{I}_{n}^{*}-\lim\chi_{A}=\chi_{T(A)}(n)$. If ${\bf x}=\sum_{i=1}^{k}c_{i}\chi_{A_{i}}$, then $Q({\bf x})_{n}=\sum_{i=1}^{k}c_{i}\chi_{T(A_{i})}(n).$ On the other hand,
$\sum_{m\in F}(\mathcal{I}_{m}^{*}-\lim{\bf x})\chi_{V_{m}}(n)$ $=(\mathcal{I}_{k}^{*}-\lim{\bf x})\chi_{V_{k}}(n)$
$=(\mathcal{I}_{n}^{*}-\lim{\bf x})\chi_{V_{n}}(n)$
$=\sum_{i=1}^{k}c_{i}\chi_{T(A_{i})}(n).$
Thus, $Q({\bf x})=\sum_{m\in F}(\mathcal{I}_{m}^{*}-\lim{\bf x})\chi_{V_{m}}$ when ${\bf x}=\sum_{i=1}^{k}c_{i}\chi_{A_{i}}$. The general case follows since the linear span of the set $\{\chi_{A}\,\colon\,A\subseteq\mathbb{N}\}$ is dense in $\ell_{\infty}$ (see [23, Proposition 2.5]). Therefore, $Q$ has the same form as given in (5.1). ∎
Remark 5.23.
We know that if $\mathcal{I}$ is an $\omega$-maximal ideal, then $\mathfrak{ad}(\mathcal{I}^{+})\leq\aleph_{0}$, as demonstrated earlier in the introduction to this section (see also [22]). The proof presented here shows also that if $\mathcal{I}$ is the intersection of a countable family of complemented ideals, then $\mathfrak{ad}(\mathcal{I}^{+})\leq\aleph_{0}$.
Now, we present examples of ideals that are not complemented. If $\mathcal{I}$ and $\mathcal{J}$ are ideals, their Fubini product ×$\mathcal{I}\times\mathcal{J}$ is the ideal on $\mathbb{N}\times\mathbb{N}$ defined by
$A\in\mathcal{I}\times\mathcal{J}\quad\mbox{if and only if}\quad\{m\in\mathbb{N}\,\colon\,\{n\in\mathbb{N}\,\colon\,(m,n)\in A\}\not\in\mathcal{J}\}\in\mathcal{I}.$
Regarding the Baire property of the Fubini product, in [11] was shown that ×$\mathcal{I}\times{\sf Fin}$ has the Baire property (hence, it is meager) for any ideal $\mathcal{I}$. On the other hand, they also showed that ${\sf Fin}\times\mathcal{I}$ has the Baire property exactly when $\mathcal{I}$ has it.
If $\mathcal{A}$ is a family of subsets of $\mathbb{N}$, the orthogonal of the family $\mathcal{A}$ is defined by
$\mathcal{A}^{\perp}=\{B\subset\mathbb{N}\,\colon\,(\forall A\in\mathcal{A})(B\cap A\in{\sf Fin})\}.$
Theorem 5.24.
Let $\mathcal{I}$ be a proper ideal on $\mathbb{N}$. Then,
  1. (1)
    $\mathfrak{ad}({\sf Fin}\times\mathcal{I})=\mathfrak{ad}(\mathcal{I}\times{\sf Fin})=2^{\aleph_{0}}$. In particular, ${\sf Fin}\times\mathcal{I}$ and ×$\mathcal{I}\times{\sf Fin}$ are not complemented.
  2. (2)
    $\mathcal{I}^{\omega\perp}$ is not complemented.
Proof.
(1) Let $\mathcal{J}\coloneqq{\sf Fin}\times\mathcal{I}$. We will prove that $\mathfrak{ad}(\mathcal{J}^{+})=2^{\aleph_{0}}$. Let $\mathcal{A}$ be a ${\sf Fin}$-AD family of size $2^{\aleph_{0}}$. For each $A\in\mathcal{A}$, define $B_{A}=A\times\mathbb{N}$. Observe that for any $m\in\mathbb{N}$, the section $\{n\in\mathbb{N}\,\colon\,(m,n)\in B_{A}\}$ is empty if $m\not\in A$, and equal to $\mathbb{N}$ if $m\in A$. Therefore, $\{B_{A}\,\colon\,A\in\mathcal{A}\}$ forms a $\mathcal{J}^{+}$-AD family of size $2^{\aleph_{0}}$. In particular, it follows that ${\sf Fin}\times\mathcal{I}$ is not complemented by Corollary 5.17.
To show that $\mathfrak{ad}(\mathcal{I}\times{\sf Fin})=2^{\aleph_{0}}$, we know that ×$\mathcal{I}\times{\sf Fin}$ has the Baire property by [11, Proposition 2.6]. Consequently, it is meager by Theorem 2.1. Therefore, we have $\mathfrak{ad}(\mathcal{I}\times{\sf Fin})=2^{\aleph_{0}}$ by Lemma 5.1. Now, from [18] it follows that ×$\mathcal{I}\times{\sf Fin}$ is not complemented.
(2) According to [23, Theorem 5.4], we have the isometry $c_{0,\mathcal{I}^{\omega\perp}}\cong c_{0}(c_{0,\mathcal{I}^{\perp}})$. By the main result of [5], the space $c_{0}(c_{0,\mathcal{I}^{\perp}})$ contains a complemented copy of $c_{0}$. Consequently, by Proposition 5.16, $c_{0,\mathcal{I}^{\omega\perp}}$ is not complemented in $\ell_{\infty}$. ∎
We say that an ideal $\mathcal{I}$ is strongly complemented if there is an operator $Q$ as in the hypothesis of Theorem 5.20. We have the following diagram summarizing some of our results:
[…]
The only examples of complemented ideals that we know are the strongly $\omega$-maximal ones. The only strict arrows we know are the ones mentioned in the diagram. To finish, we present the examples showing that the bottom arrows are strict. Example 3.8 shows that the first arrow is strict. For the second arrow, an example is given by [15, Lemma 2.2, Corollary 4.1]. Finally, for the last arrow, we mentioned above ${\sf Fin}\times\mathcal{I}$ is not meager whenever $\mathcal{I}$ is a maximal ideal, but $\mathfrak{ad}({\sf Fin}\times\mathcal{I})=2^{\aleph_{0}}$.
Acknowledgments
We would like to thank Alan Dow for suggesting the Example 3.8. We also thank the partial support given by the VIE-grant 4248 of Universidad Industrial de Santander.

### v2

@@
 Abstract.
-We study the complementation (in $\ell_{\infty}$) of the Banach space $c_{0,\mathcal{I}}$, consisting of all bounded sequences $(x_{n})$ that $\mathcal{I}$-converge to 00, endowed with the supremum norm, where $\mathcal{I}$ is an ideal of subsets of $\mathbb{N}$.
+We study the complementation (in $\ell_{\infty}$) of the Banach space $c_{0,\mathcal{I}}$, consisting of all bounded sequences $(x_{n})$ that $\mathcal{I}$-converge to 0, endowed with the supremum norm, where $\mathcal{I}$ is an ideal of subsets of $\mathbb{N}$.
 We show that the complementation of these spaces is related to a condition requiring that the ideal is the intersection of a countable family of maximal ideals, which we refer to as $\omega$-maximal ideals.
-We prove that if $c_{0,\mathcal{I}}$ admits a projection satisfying a certain condition, then $\mathcal{I}$ must be a special type of $\omega$-maximal ideal.
+We prove that $\mathcal{I}$ is $\omega$-maximal exactly when $c_{0,\mathcal{I}}$ is the kernel of an operator from $\ell_{\infty}$ to itself satisfying certain property.
+In addition, we show that the existence of a Banach lattice embedding from $\ell_{\infty}/c_{0,\mathcal{I}}$ into $\ell_{\infty}$ is equivalent to $\mathcal{I}$ being a $\omega$-maximal ideal.
+Finally, we show that $\mathcal{I}$ is a strongly $\omega$-maximal ideal if and only if the quotient $\ell_{\infty}/c_{0,\mathcal{I}}$ is Banach lattice isomorphic to $\ell_{\infty}$.
 Additionally, we characterize when the quotient space $c_{0,\mathcal{J}}/c_{0,\mathcal{I}}$ is finite-dimensional for two ideals $\mathcal{I}\subsetneq\mathcal{J}$.
@@
 We call a proper ideal $\mathcal{I}$ complemented if $c_{0,\mathcal{I}}$ is complemented in $\ell_{\infty}$.
+There has been some recent interest in this notion [13, 15, 20].
 Leonetti [20] proved that any meager ideal is not complemented, where a meager ideal refers to an ideal that is meager as a subset of the Cantor cube $\{0,1\}^{\mathbb{N}}$, identified via characteristic functions.
@@
 We extend Kania’s result to certain special $\omega$-maximal ideals.
+Finally, continuing this approach, Hrušák-Sáenz [13] found very recently a characterization of complemented ideals in terms of the quotient $\ell_{\infty}/c_{0,\mathcal{I}}$.
+Namely, $\mathcal{I}$ is complemented if and only if the quotient $\ell_{\infty}/c_{0,\mathcal{I}}$ is isomorphic to a subspace of $\ell_{\infty}$.
+Motivated by the discussion above, our work focuses on analyzing the structural properties of complemented ideals.
 We say that an ideal is strongly $\omega$-maximal if there exists a collection $\{\mathcal{I}_{n}\colon n\in\mathbb{N}\}$ of maximal ideals such that $\mathcal{I}=\bigcap_{n}\mathcal{I}_{n}$, and the family $\{\mathcal{I}_{n}^{*}\colon n\in\mathbb{N}\}$ is discrete in $\beta\mathbb{N}$ (where $\mathcal{I}^{*}$ denotes the dual filter and $\beta\mathbb{N}$ is the Stone–Čech compactification of $\mathbb{N}$).
@@
 In particular, we obtain that $\ell_{\infty}/c_{0,\mathcal{I}}$ is finite-dimensional if and only if $\mathcal{I}$ is a finite intersection of maximal ideals.
-In the last section, we discuss the problem of the complementation of $c_{0,\mathcal{I}}$ in $\ell_{\infty}$.
-One of our main results is that any strongly $\omega$-maximal ideal is complemented (see Theorem 5.8).
-We provide two different proofs of this result.
-The first one (see Theorem 5.2) is shorter but considerably less informative, as it relies on a classical theorem of Lindenstrauss stating that a closed subspace of $\ell_{\infty}$ is complemented if and only if it is isomorphic to $\ell_{\infty}$.
+In Section 5, we discuss properties of complemented ideals.
+We show that the sum and the countable intersection of complemented ideals is complemented.
+So, in particular, every $\omega$-maximal ideal is complemented.
+We also prove that any strongly $\omega$-maximal ideal is complemented by showing an explicit projection (see Theorem 5.14).
 Furthermore, we characterize strongly $\omega$-maximal ideals in terms of the type of projections on their associated space $c_{0,\mathcal{I}}$ (see Theorem 5.16).
-In addition, we show that if $c_{0,\mathcal{I}}$ is complemented in $\ell_{\infty}$ by a projection satisfying an extra condition, then $\mathcal{I}$ must be an $\omega$-maximal ideal (see Theorem 5.18).
-Whether the converse holds—that is, whether every $\omega$-maximal ideal is complemented—remains an open question.
+In addition, we show that the existence of a Banach lattice embedding from $\ell_{\infty}/c_{0,\mathcal{I}}$ into $\ell_{\infty}$ to $\ell_{\infty}$ is equivalent to $\mathcal{I}$ being a $\omega$-maximal ideal (see Theorem 5.11).
+Finally, we show that $\mathcal{I}$ is a strongly $\omega$-maximal ideal if and only if the quotient $\ell_{\infty}/c_{0,\mathcal{I}}$ is Banach lattice isomorphic to $\ell_{\infty}$ (see Theorem 5.19).
 We show that $c_{0}$ is not complemented in $c_{0,\mathcal{J}}$ for any ideal $\mathcal{J}$ properly extending $\mathsf{Fin}$.
@@
 All Banach lattices analyzed here are assumed to be real.
-However, our results can be extended to complex Banach lattices in the usual manner [24, Chapter 2, p.
+However, our results can be extended to complex Banach lattices in the usual manner [26, Chapter 2, p.
 133].
@@
 In addition, if $E$ and $X$ are Banach lattices, $P$ is called positive if $Px\geq{\bf 0}$ for all $x\geq{\bf 0}$.
+Also, a bounded linear operator $T\colon E\to X$ is called Banach lattice homomorphism if $T(x\vee y)=Tx\vee Ty$ for each $x,y\in E$.
 An ideal $\mathcal{I}$ on a set $X$ is a collection of subsets of $X$ satisfying:
@@
 Since we are always assuming that $\sf Fin$} is contained in any ideal under consideration, every maximal ideal extending a given ideal is necessarily non-principal.
-It is worth keeping in mind that the intersection of less than $2^{\aleph_{0}}$ maximal ideals on $\mathbb{N}$ does not have the Baire property, and hence it is non-meager ([22], [26, Proposition 23]).
+It is worth keeping in mind that the intersection of less than $2^{\aleph_{0}}$ maximal ideals on $\mathbb{N}$ does not have the Baire property, and hence it is non-meager ([24], [28, Proposition 23]).
 Our main interest will be on $\omega$-maximal ideals since they are related to the complementation of $c_{0,\mathcal{I}}$ in $\ell_{\infty}$.
@@
   2. (2)
-[…]
+    $\mathcal{I}$ is a limit point of $\mathfrak{D}$ if and only if $\bigcap\mathfrak{D}=\bigcap(\mathfrak{D}\setminus\{\mathcal{I}\})$.
+Consequently, $\mathfrak{D}$ is discrete if and only if $\bigcap(\mathfrak{D}\setminus\{\mathcal{K}\})\not\subseteq\mathcal{K}$ for each $\mathcal{K}\in\mathfrak{D}$.
 Proof.
@@
     It follows from (1).∎
-For each collection of maximal ideals $\mathfrak{E}=\{\mathcal{I}_{n}\,\colon\,n\in\mathbb{N}\}$ we define a family of infinite subsets of $\mathbb{N}$ which will be very helpfull for what follows:
+For each collection of maximal ideals $\mathfrak{E}=\{\mathcal{I}_{n}\,\colon\,n\in\mathbb{N}\}$ we define a family of infinite subsets of $\mathbb{N}$ which will be very helpful for what follows:
 $\mathcal{DI}(\mathfrak{E})=\{M\in[\mathbb{N}]^{\omega}\,\colon\,\mbox{$\{\mathcal{I}_{n}\,\colon\,n\in M\}$ is discrete}\}.$
@@
 Let $L=\{n\in\mathbb{N}:J_{n}\in\mathfrak{E}_{2}\}$ and notice that $L\in\mathcal{DI}(\mathfrak{E}_{1}\cup\mathfrak{E}_{2})$ is maximal by Lemma 3.3 and thus $n_{0}\in M$.
-On the other hand, $\bigcap\{J_{n}:n\in M\setminus\{n_{0}\}\}=\bigcap\mathfrak{E}_{2}\subseteq J_{n_{0}}$ and, by Lemma 3.2, $M$ is not discrete, a contradiction.
+On the other hand, $\bigcap\{J_{n}:n\in M\setminus\{n_{0}\}\}=\bigcap\mathfrak{E}_{2}\subseteq J_{n_{0}}$ and, by Proposition 3.2, $M$ is not discrete, a contradiction.
 ∎
@@
 Suppose $\mathfrak{E}^{*}$ is scattered.
-Since the collection of isolated points of $\mathfrak{E}^{*}$ is discrete and dense in $\mathfrak{E}^{*}$ (see [25, p.
+Since the collection of isolated points of $\mathfrak{E}^{*}$ is discrete and dense in $\mathfrak{E}^{*}$ (see [27, p.
 150]), there is $M\in\mathcal{DI}(\mathfrak{E})$ such that $\{\mathcal{I}^{*}:\,n\in M\}$ is dense in $\mathfrak{E}^{*}$, then easily $M$ is maximal in $\mathcal{DI}(\mathfrak{E})$.
 Thus by Theorem 3.4, $\mathcal{I}$ is strongly $\omega$-maximal.
-The converse is obvious.
+The converse obviously follows from Theorem 3.4.
 ∎
 Remark 3.6.
-[…]
+Following [16], a collection of non principal ultrafilters $\mathfrak{M}$ is said minimal, if $\bigcap\mathfrak{M}\neq\bigcap(\mathfrak{M}\setminus\{\mathcal{F}\})$ for every $\mathcal{F}\in\mathfrak{M}$.
+A filter $\mathcal{F}$ has a minimal representation if $\mathcal{F}=\bigcap\mathfrak{M}$ for some minimal collection $\mathfrak{M}$ of ultrafilters.
+Lemma 3.2 says that $\mathfrak{M}$ is minimal if and only if it is a discrete subset of $\beta\mathbb{N}$.
+And, from Theorem 3.5 it follows that a filter $\mathcal{F}$ has a countable minimal representation if and only if $\mathcal{F}^{*}$ is strongly $\omega$-maximal ideal.
+They also showed that whenever a filter admits a countable minimal representation, it is unique.
+This also follows from Theorem 3.4.
 We present two examples.
@@
 Since this space is zero-dimensional and has no isolated points, for every discrete $D\subseteq{\mathbb{N}}^{<\omega}$ there is $s\in{\mathbb{N}}^{<\omega}\setminus D$ such that $D\cup\{s\}$ is still discrete, i.e., there are no maximal discrete subsets of ${\mathbb{N}}^{<\omega}$.
-It is a classical fact that every countable extremally disconnected Hausdorff space can be embedded in $\beta\mathbb{N}$ (see, for instance, [27, Theorem 1.4.7]).
-Therefore, if $\mathfrak{E}\subseteq\beta\mathbb{N}$ is homeomorphic to our space, then $\mathfrak{E}$ has no maximal discrete subsets, that is =$\mathcal{I}=\bigcap\{\mathcal{U}^{*}:\mathcal{U}\in\mathfrak{E}\}$ is an $\omega$-maximal ideal which is not strongly $\omega$-maximal (by Theorem 3.4).
+It is a classical fact that every countable extremally disconnected Hausdorff space can be embedded in $\beta\mathbb{N}$ (see, for instance, [29, Theorem 1.4.7]).
+Therefore, if $\mathfrak{E}\subseteq\beta\mathbb{N}$ is homeomorphic to our space, then $\mathfrak{E}$ has no maximal discrete subsets, that is $\mathcal{I}=\bigcap\{\mathcal{U}^{*}:\mathcal{U}\in\mathfrak{E}\}$ is an $\omega$-maximal ideal which is not strongly $\omega$-maximal (by Theorem 3.4).
 Our next result implies that the $\omega$-maximal ideal constructed in the previous example is nowhere maximal.
@@
 Conversely, suppose $\mathcal{I}$ is $k$-maximal in $\mathcal{J}$ and let $\{\mathcal{L}_{1},\ldots,\mathcal{L}_{k}\}$ be a family of maximal ideals on $\mathbb{N}$ with $\mathcal{J}\not\subset\mathcal{L}_{i}$ for all $1\leq i\leq k$ and $\mathcal{I}=(\bigcap_{i=1}^{k}\mathcal{L}_{i})\cap\mathcal{J}$.
-By Theorem 3.1, there is a partition $\{D_{1},\ldots,D_{k}\}$ of $\mathbb{N}$ such that $D_{i}\in\mathcal{L}^{*}_{i}\cap\mathcal{L}_{j}$ for all $j,i$ in $\{1,\ldots,k\}$ with $j\neq i$.
+By Proposition 3.1, there is a partition $\{D_{1},\ldots,D_{k}\}$ of $\mathbb{N}$ such that $D_{i}\in\mathcal{L}^{*}_{i}\cap\mathcal{L}_{j}$ for all $j,i$ in $\{1,\ldots,k\}$ with $j\neq i$.
 For each $1\leq i\leq k$, let $X_{i}\in\mathcal{J}\setminus\mathcal{L}_{i}$.
@@
 Proof.
-[…]
+Let $A_{1},\ldots,A_{m}\in\mathcal{A}$ be such that $\sum_{j=1}^{m}a_{j}(\chi_{A_{j}}+c_{0,\mathcal{I}})={\bf 0}$ for some $\{a_{j}\,\colon\,1\leq j\leq m\}\subset\mathbb{K}$.
+Let $D_{1}=A_{1}$ and $D_{j}=A_{j}\setminus(A_{1}\cup\cdots\cup A_{j-1})$ for $2\leq j\leq m$.
+Since $A_{i}\cap A_{j}\in\mathcal{I}$ for all $i,j\in\{1,\ldots,k\}$ with $i\neq j$, we have $D_{j}\in\mathcal{J}\cap\mathcal{I}^{+}$ and $\chi_{A_{j}}+c_{0,\mathcal{I}}=\chi_{D_{j}}+c_{0,\mathcal{I}}$ for all $j\in\{1,\ldots,m\}$.
+So, $\sum_{j=1}^{m}a_{j}(\chi_{D_{j}}+c_{0,\mathcal{I}})={\bf 0}$.
+Thus, $a_{j}=0$ for all $1\leq j\leq m$.
+∎
 Theorem 4.6.
@@
 Proof.
-[…]
+Suppose that $\dim(c_{0,\mathcal{J}}/c_{0,\mathcal{I}})=k$.
+By [26, Corollary 1, p.
+70], there is an order isomorphism $\phi\colon c_{0,\mathcal{J}}/c_{0,\mathcal{I}}\to\mathbb{R}^{k}$.
+Let $\Lambda\colon c_{0,\mathcal{J}}\to\mathbb{R}^{k}$ be given by $\Lambda({\bf y})=\phi({\bf y}+c_{0,\mathcal{I}})$ if ${\bf y}\in c_{0,\mathcal{J}}$.
+Notice that $\Lambda$ is an onto Banach lattice homomorphism.
+For each $j\in\{1,\ldots,k\}$, let ${\bf y}_{j}\in c_{0,\mathcal{J}}\setminus c_{0,\mathcal{I}}$ be such that $\Lambda({\bf y}_{j})=e_{j}$.
+Fix $j\in\{1,\ldots,k\}$ and let $\varepsilon_{j}>0$ be such that $A_{j}:=A(\varepsilon_{j},{\bf y}_{j})\not\in\mathcal{I}$.
+So, $\varepsilon_{j}\chi_{A_{j}}\leq{\bf y}_{j}$.
+Since $\Lambda$ is order preserving, we conclude that $\Lambda(\chi_{A_{j}})=a_{j}e_{j}$ for some $a_{j}>0$.
+Notice that if $i,j\in\{1,\ldots,k\}$ and $i\neq j$, we have $A_{i}\cap A_{j}\in\mathcal{I}$ because of $\Lambda(\chi_{A_{i}\cap A_{j}})=\Lambda(\chi_{A_{i}}\wedge\chi_{A_{j}})=a_{i}e_{i}\wedge a_{j}e_{j}={\bf 0}$.
+Thus, $\{A_{1},\ldots,A_{k}\}$ is a $\mathcal{I}$-AD family.
+As in the proof of Lemma 4.5, we will assume that the family is pairwise disjoint.
+From Lemma 4.5 we obtain that $\{\chi_{A_{j}}+c_{0,\mathcal{I}}\,\colon\,1\leq j\leq k\}$ is a basis for $c_{0,\mathcal{J}}/c_{0,\mathcal{I}}$.
+We claim that $\{A_{1},\ldots,A_{k}\}$ is a $(\mathcal{I},\mathcal{J},k)$-family:
   1. (1)
@@
   2. (2)
-    Since $A_{i}\cap A_{j}\in\mathcal{I}$ for all $i,j\in\{1,\ldots,k\}$ and $i\neq j$, by a standard procedure we can make them pairwise disjoint and thus we may assume that $A_{i}\cap A_{j}=\emptyset$ for each $i,j\in\{1,\ldots,k\}$ with $i\neq j$.
+    $\{A_{j}:1\leq j\leq k\}$ is pairwise disjoint by construction.
   3. (3)
@@
 Then $\dim(\ell_{\infty}/c_{0,\mathcal{I}})=k$ if and only if $\mathcal{I}$ is the intersection of exactly $k$ maximal ideals on $\mathbb{N}$.
-5. When is $c_{0,\mathcal{I}}$ complemented in $\ell_{\infty}$?
-In this section we address the problem of when an ideal $\mathcal{I}$ is complemented, that is, when $c_{0,\mathcal{I}}$ is complemented in $\ell_{\infty}$.
-We recall that a classical theorem of Lindenstrauss [19], which states that a closed subspace $X\subseteq\ell_{\infty}$ is complemented if and only if $X$ is isomorphic to $\ell_{\infty}$.
-As it was noticed by Kania in [14], if $\mathcal{I}$ is the intersection of finitely many maximal ideals, then $\mathcal{I}$ is complemented (for a proof see [23, Proposisition 5.20]).
-Therefore, if $\mathcal{I}$ is the intersection of finitely many maximal ideals on an infinite set $A$, then $c_{0,\mathcal{I}}$ is isomorphic to $\ell_{\infty}(A)$ (and hence, isomorphic to $\ell_{\infty}$).
-Recall that a $\mathcal{A}\subseteq\mathcal{I}^{+}$ is $\mathcal{I}$-AD if $X\cap Y\in\mathcal{I}$ for all $X,Y\in\mathcal{A}$. We let
-[…]
-We will write just $\mathfrak{ad}(\mathcal{A})$ when it is clear from the context which ideal $\mathcal{I}$ is used.
-We begin by stating a lemma which is a well-known consequence of Theorem 2.1 (for a proof, see, for instance, [18]).
-Lemma 5.1.
-Let $\mathcal{I}$ be a meager ideal on $\mathbb{N}$. Then $\mathfrak{ad}(\mathcal{\mathcal{I}}^{+})=2^{\aleph_{0}}$.
-On the other hand, $\mathfrak{ad}(\mathcal{\mathcal{I}}^{+})\leq\aleph_{0}$ for every $\omega$-maximal ideal $\mathcal{I}$, this was implicitly shown by Plewik [22] and, for the sake of completion, we include a direct proof.
-Let $\mathcal{I}=\bigcap_{n}\mathcal{I}_{n}$ be an $\omega$-maximal ideal and suppose $\mathcal{B}\subseteq\mathcal{I}^{+}$ is an uncountable $\mathcal{I}$-AD family.
-Then, there is $n$ such that $|\mathcal{B}\cap\mathcal{I}_{n}^{+}|\geq 2$.
-If $A,B\in\mathcal{B}\cap\mathcal{I}_{n}^{+}$ are two different sets, then $A\cap B\in\mathcal{I}\subseteq\mathcal{I}_{n}$, which contradicts that $\mathcal{I}_{n}$ is maximal.
-Additionally, any $\omega$-maximal ideal is not measurable with respect to the usual product measure on $2^{\mathbb{N}}$ (see [15, section 4.1]).
-[…]
+5. Structural properties of complemented ideals
+In this section, we study the structural properties of complemented ideals $\mathcal{I}$, that is, ideals for which the Banach space $c_{0,\mathcal{I}}$ is complemented in $\ell_{\infty}$.
+We recall that a classical theorem of Lindenstrauss [21] says that a closed subspace $X\subseteq\ell_{\infty}$ is complemented if and only if $X$ is isomorphic to $\ell_{\infty}$.
+As noted by Kania in [15], the intersection of finitely many maximal ideals is complemented (for a proof, see [25, Proposisition 5.20]).
+Therefore, if $\mathcal{I}$ is such an ideal, then $c_{0,\mathcal{I}}$ is isomorphic to $\ell_{\infty}$.
+Recently, Hrušák and Sáenz characterized the complementation of $c_{0,\mathcal{I}}$ in terms of the quotient $\ell_{\infty}/c_{0,\mathcal{I}}$ [13].
+Before stating their result, we need to introduce some notation.
+For an ideal $\mathcal{I}$, let $K_{\mathcal{I}}=\{p\in\beta\mathbb{N}\,\colon\,\mathcal{I}^{*}\subseteq p\}$.
+It is known that $K_{\mathcal{I}}$ is a compact Hausdorff space, since it is a closed subset of $\beta\mathbb{N}$.
+Moreover, the map
+$\Psi\colon\ell_{\infty}/c_{0,\mathcal{I}}$ $\to C(K_{\mathcal{I}})$
+${\bf x}+c_{0,\mathcal{I}}$ $\mapsto\Psi({\bf x}+c_{0\mathcal{I}})\,\colon\,p\in K_{\mathcal{I}}\mapsto p-\lim{\bf x}$
+is a Banach lattice isometry. Therefore, we have the following:
+Theorem 5.1.
+(Kania [15]) $\ell_{\infty}/c_{0,\mathcal{I}}$ is Banach lattice isometric to $C(K_{\mathcal{I}})$.
+Theorem 5.2.
+(Hrušák and Sáenz [13]). Let $\mathcal{I}$ be an ideal on $\mathbb{N}$. Then, the following statements are equivalent:
+  1. (1)
+    $\mathcal{I}$ is complemented.
+  2. (2)
+    $\ell_{\infty}/c_{0,\mathcal{I}}$ is isomorphic to a subspace of $\ell_{\infty}$.
+  3. (3)
+    $\ell_{\infty}/c_{0,\mathcal{I}}$ is isometric to a subspace of $\ell_{\infty}$.
+  4. (4)
+    $C(K_{\mathcal{I}})$ is isomorphic to a subspace of $\ell_{\infty}$.
+Corollary 5.3.
+Let $\mathcal{I}$ be an ideal on $\mathbb{N}$. Then, $\mathcal{I}$ is complemented if and only if $c_{0,\mathcal{I}}$ is the kernel of a continuous operator from $\ell_{\infty}$ to $\ell_{\infty}$.
+Proof.
+If $\mathcal{I}$ is complemented, then $c_{0,\mathcal{I}}$ is a kernel of an operator from $\ell_{\infty}$ to $\ell_{\infty}$.
+Conversely, let $Q\colon\ell_{\infty}\to\ell_{\infty}$ be a continuous operator such that $\ker Q=c_{0,\mathcal{I}}$.
+Then, the map $\hat{Q}\colon\ell_{\infty}/c_{0,\mathcal{I}}\to\ell_{\infty}$ defined by ${\bf x}+c_{0,\mathcal{I}}\mapsto Q({\bf x})$, is an embedding.
+By Theorem 5.2, $\mathcal{I}$ is complemented.
+∎
+Recall that if $(X_{j})_{j\in\mathbb{N}}$ is a sequence of Banach spaces, the space $\ell_{\infty}((X_{j})_{j\in\mathbb{N}})$ denotes their $\ell_{\infty}$-sum, i.e., the Banach space of all bounded sequences $(x_{j})\in\prod_{j}X_{j}$, equipped with the norm $\|\cdot\|$ given by $\|(x_{j})\|=\sup_{j}\|x_{j}\|$.
+Remark 5.4.
+Let $(X_{n})_{n\in\mathbb{N}}$ and $(Y_{n})_{n\in\mathbb{N}}$ be sequences of Banach spaces (lattices) and assume that $T_{n}\colon X_{n}\to Y_{n}$ is an isomorphism (Banach lattice isomorphism) for each $n\in\mathbb{N}$.
+Then, $\ell_{\infty}((X_{n})_{n\in\mathbb{N}})$ is isomorphic (Banach lattice isomorphic) to $\ell_{\infty}((Y_{n})_{n\in\mathbb{N}})$ via the map $T\colon(x_{n})\mapsto(T_{n}(x_{n}))$.
+In particular, if $\{K_{n}\,\colon\,n\in\mathbb{N}\}$ is a partition of $\mathbb{N}$, then $\ell_{\infty}((\ell_{\infty}(K_{m}))_{m\in\mathbb{N}})$ is isomorphic to $\ell_{\infty}$, since $\ell_{\infty}(K_{n})\sim\ell_{\infty}$ for each $n\in\mathbb{N}$ and $\ell_{\infty}(\ell_{\infty})\sim\ell_{\infty}$.
 Our first result about complemented ideals is based on Lindenstrauss’s theorem and a result from [25].
@@
 Then, the ideal $\mathcal{J}=\bigoplus_{n\in\mathbb{N}}\mathcal{I}_{n}$ is complemented.
-In particular, if $\mathcal{I}$ is a strongly $\omega$-maximal ideal, then $\mathcal{I}$ is complemented.
 Proof.
-By the aforementioned Lindenstrauss’s theorem, $c_{0,\mathcal{I}_{m}}$ is isomorphic to $\ell_{\infty}(K_{m})$ for each $m\in\mathbb{N}$. From [23, Theorem 5.2] it follows that
+By the aforementioned Lindenstrauss’s theorem, $c_{0,\mathcal{I}_{m}}$ is isomorphic to $\ell_{\infty}(K_{m})$ for each $m\in\mathbb{N}$. From [25, Theorem 5.2] and Remark 5.4 it follows that
 $c_{0,\mathcal{J}}\cong\ell_{\infty}((c_{0,\mathcal{I}_{m}})_{m\in\mathbb{N}})\sim\ell_{\infty}((\ell_{\infty}(K_{m}))_{m\in\mathbb{N}})\sim\ell_{\infty}.$
-Again, by Lindenstrauss’s theorem, we conclude that $c_{0,\mathcal{J}}$ is complemented in $\ell_{\infty}$.
-Suppose $\mathcal{I}$ is strongly $\omega$-maximal.
-By Proposition 3.1, there is a partition $\{K_{n}\}$ of $\mathbb{N}$ such that $\mathcal{I}=\bigoplus_{n\in\mathbb{N}}(\mathcal{I}_{n}\restriction K_{n})$ and $\mathcal{I}_{n}\restriction K_{n}$ is maximal on $K_{n}$ for each $n$.
-Then, the previous argument shows that $c_{0,\mathcal{I}}$ is isomorphic to $\ell_{\infty}$.
-∎
+Again, by Lindenstrauss’s theorem, we conclude that $c_{0,\mathcal{J}}$ is complemented in $\ell_{\infty}$. ∎
 Corollary 5.6.
@@
 Conversely, suppose that $\mathcal{I}^{\omega}$ is complemented.
-By [23, Theorem 5.2] we have $c_{0,\mathcal{I}^{\omega}}\cong\ell_{\infty}(c_{0,\mathcal{I}})$.
+By [25, Theorem 5.2] we have $c_{0,\mathcal{I}^{\omega}}\cong\ell_{\infty}(c_{0,\mathcal{I}})$.
 Consequently, $c_{0,\mathcal{I}}$ is isomorphic to a complemented subspace of $c_{0,\mathcal{I}^{\omega}}$.
-On the other hand, notice that $c_{0,\mathcal{I}^{\omega}}\sim\ell_{\infty}$ by Lindenstrauss’s theorem.
-Thus, $c_{0,\mathcal{I}}$ is isomorphic to a complemented subspace of $\ell_{\infty}$.
-Therefore, $\mathcal{I}$ is complemented.
+On the other hand, by Lindenstrauss’s theorem, $c_{0,\mathcal{I}^{\omega}}\sim\ell_{\infty}$.
+Hence, $c_{0,\mathcal{I}}$ is isomorphic to a complemented subspace of $\ell_{\infty}$.
+Therefore, by Lindenstrauss’s theorem, $c_{0,\mathcal{I}}$ is isomorphic to $\ell_{\infty}$, and thus $\mathcal{I}$ is complemented.
 ∎
-Proposition 5.4.
-Let $\mathcal{I}$ and $\mathcal{J}$ be ideals on $\mathbb{N}$ with $\mathcal{I}\subsetneq\mathcal{J}$ and $k$ be a positive integer.
-If $\mathcal{J}$ is complemented and $\mathcal{I}$ is $k$-maximal in $\mathcal{J}$, then $\mathcal{I}$ is complemented.
+Theorem 5.7.
+Let $\{\mathcal{I}_{n}\,\colon\,n\in\mathbb{N}\}$ be a collection of complemented ideals.
+Then, $\mathcal{I}=\bigcap_{n\in\mathbb{N}}\mathcal{I}_{n}$ is complemented.
+In particular, every $\omega$-maximal ideal is complemented.
 Proof.
-Since $\mathcal{I}$ is $k$-maximal in $\mathcal{J}$, it follows that $c_{0,\mathcal{I}}$ is complemented in $c_{0,\mathcal{J}}$ by Theorem 4.6.
-Additionally, $c_{0,\mathcal{J}}$ is complemented in $\ell_{\infty}$.
-Therefore, $c_{0,\mathcal{I}}$ is complemented in $\ell_{\infty}$.
-This completes the proof.
+Let $A\colon\ell_{\infty}\to\ell_{\infty}((\ell_{\infty}/c_{0,\mathcal{I}_{n}})_{n\in\mathbb{N}})$ be defined by $A({\bf x})=({\bf x}+c_{0,\mathcal{I}_{n}})$, $\bf x$$\in\ell_{\infty}$.
+Then $A$ is linear and continuous, and $\ker A=c_{0,\mathcal{I}}$.
+Hence, $\ell_{\infty}/c_{0,\mathcal{I}}$ is isomorphic to a subspace of $\ell_{\infty}((\ell_{\infty}/c_{0,\mathcal{I}_{n}})_{n\in\mathbb{N}})$.
+On the other hand, since $\mathcal{I}_{n}$ is complemented for each $n\in\mathbb{N}$, it follows from Theorem 5.2, that $\ell_{\infty}/c_{0,\mathcal{I}_{n}}$ is isomorphic to $\ell_{\infty}$ for every $n\in\mathbb{N}$.
+Therefore, $\ell_{\infty}((\ell_{\infty}/c_{0,\mathcal{I}_{n}})_{n\in\mathbb{N}})\sim\ell_{\infty}(\ell_{\infty})\sim\ell_{\infty}$.
+Consequently, $\ell_{\infty}/c_{0,\mathcal{I}}$ is isomorphic to a subspace of $\ell_{\infty}$ and $\mathcal{I}$ is complemented.
 ∎
 Remark 5.8.
-Despite these results, it remains unknown whether the intersection of a countable family of complemented ideals is itself complemented.
-Even in the case of two complemented ideals, $\mathcal{I}$ and $\mathcal{J}$, we do not know whether $\mathcal{I}\cap\mathcal{J}$ is complemented.
+Example 2 in [13] shows that the second part of Theorem 5.7 is not an equivalence.
 We say that a proper ideal $\mathcal{I}$ on a set $A\subseteq\mathbb{N}$ is complemented in $A$ if $c_{0,\mathcal{I}}$ is complemented in $\ell_{\infty}(A)$ (recall that $c_{0,\mathcal{I}}$ consists of sequences of the form $(x_{n})_{n\in A}$).
@@
 Proposition 5.9.
-Let $\mathcal{I}$ be an ideal on $\mathbb{N}$ and let $A$ and $B$ be infinite subsets of $\mathbb{N}$ with $A\subseteq B$ and $A\in\mathcal{I}^{+}$.
+Let $\mathcal{I}$ be an ideal on $\mathbb{N}$, and let $A$ and $B$ be infinite subsets of $\mathbb{N}$ with $A\subseteq B$ and $A\in\mathcal{I}^{+}$.
 If $\mathcal{I}\restriction B$ is complemented in $B$, then $\mathcal{I}\restriction A$ is complemented in $A$.
@@
 defines a projection, completing the proof. ∎
-Now we present a more informative proof of Theorem 5.2.
-But first we need an auxiliary result.
+We will show that, under an additional condition, the complementation of $\mathcal{I}$ is equivalent to $\mathcal{I}$ being $\omega$-maximal.
+For that end, we need a key lemma whose argument is motivated by the proof of the non-complementation of $c_{0}$ in $\ell_{\infty}$ due to Whitley (see [7, Theorem 5.6] or [1, Theorem 2.5.4]).
+Lemma 5.10.
+Let $\mathcal{I}\subsetneq\mathcal{J}$ be ideals on $\mathbb{N}$.
+Suppose that there is a linear bounded operator $Q\colon c_{0,\mathcal{J}}\to c_{0,\mathcal{J}}$ such that $\ker Q=c_{0,\mathcal{I}}$.
+Then, there exists a collection $\{\mathcal{A}_{n,k}\colon\,n,k\in\mathbb{N}\}$ of subsets of $2^{\mathbb{N}}$ satisfying:
+  1. (1)
+    $\mathcal{A}_{n,k}\subseteq\mathcal{J}\cap\mathcal{I}^{+}$ for each $n,k\in\mathbb{N}$.
+  2. (2)
+    $\mathcal{A}_{n,k}\subseteq\mathcal{A}_{n,k+1}$ for all $n,k\in\mathbb{N}$.
+  3. (3)
+    There is $M>0$ such that $\mathfrak{ad}_{\mathcal{I}}(\mathcal{A}_{n,k})\leq kM$ for all $n,k\in\mathbb{N}$.
+  4. (4)
+    For each $n\in\mathbb{N}$, let $\mathcal{A}_{n}=\bigcup_{k\in\mathbb{N}}\mathcal{A}_{n,k}$. Then $\mathcal{J}\cap\mathcal{I}^{+}=\bigcup_{n\in\mathbb{N}}\mathcal{A}_{n}$.
+  5. (5)
+    If $A\in\mathcal{A}_{n,k}$ and $A=B\cup C$ with $B\cap C\not\in\mathcal{A}_{n}$, then either $B\in\mathcal{A}_{n,2k}$ or $C\in\mathcal{A}_{n,2k}$.
+In particular, if $A\in\mathcal{A}_{n}$ and $A=B\cup C$ with $B\cap C\not\in\mathcal{A}_{n}$, then either $B\in\mathcal{A}_{n}$ or $C\in\mathcal{A}_{n}$.
+Proof.
+For each $n,k\in\mathbb{N}$, define
+$\mathcal{A}_{n,k}=\{A\in\mathcal{J}\,\colon\,|Q(\chi_{A})_{n}|\geq 1/k\}.$
+Since $\mathcal{I}=\{A\in\mathcal{J}\,\colon\,Q(\chi_{A})={\bf 0}\}$, we have
+$\mathcal{J}\cap\mathcal{I}^{+}$ $=\{A\in\mathcal{J}\,\colon\,Q(\chi_{A})\neq{\bf 0}\}$
+$=\bigcup_{k,n\in\mathbb{N}}\{A\in\mathcal{J}\,\colon\,|Q(\chi_{A})_{n}|\geq 1/k\}$
+$=\bigcup_{n\in\mathbb{N}}\mathcal{A}_{n}.$
+Then (1) and (4) hold.
+(2) is obvious.
+To see (3), let $m\in\mathbb{N}$ be given and $A_{1},\ldots,A_{m}\in\mathcal{A}_{n,k}$ be such that $A_{i}\cap A_{j}\in\mathcal{I}$ for $i\neq j$.
+If $F_{1}=A_{1}$ and $F_{j}=A_{j}\setminus(A_{1}\cup\cdots\cup A_{j-1})$ for $2\leq j\leq m$, then $Q(\chi_{A_{j}})=Q(\chi_{F_{j}})$ for all $1\leq j\leq m$.
+For $1\leq j\leq m$, let $a_{j}\in\mathbb{K}$ be such that $a_{j}Q(\chi_{F_{j}})_{n}=|Q(\chi_{F_{j}})_{n}|$.
+Thus,
+$\frac{m}{k}\leq\sum_{j=1}^{m}|Q(\chi_{A_{j}})_{n}|$ $=\sum_{j=1}^{m}|Q(\chi_{F_{j}})_{n}|$
+$=\sum_{j=1}^{m}a_{j}Q(\chi_{F_{j}})_{n}$
+$\leq\left\|\sum_{j=1}^{m}a_{j}Q(\chi_{F_{j}})\right\|$
+$\leq\|Q\|\left\|\sum_{j=1}^{m}a_{j}\chi_{F_{j}}\right\|=\|Q\|.$
+Hence, $m\leq k\|Q\|$.
+Finally, we show (5).
+Let $A\in\mathcal{A}_{n,k}$ and $A=B\cup C$ with $B\cap C\not\in\mathcal{A}_{n}$.
+Then $|Q(\chi_{B\cap C})_{n}|=0$.
+Assume that $|Q(\chi_{B})_{n}|<1/2k$ and $|Q(\chi_{C})_{n}|<1/2k$.
+Note that $Q(\chi_{B})_{n}=Q(\chi_{B\setminus C})_{n}$ and $Q(\chi_{C})_{n}=Q(\chi_{C\setminus B})_{n}$.
+Thus
+$|Q(\chi_{A})_{n}|=|Q(\chi_{A\setminus B\cap C})_{n}|=|Q(\chi_{B})_{n}+Q(\chi_{C})_{n}|<1/k,$
+which is absurd. ∎
+Now we prove, with an extra assumption, that the complementation of $c_{0,\mathcal{I}}$ is equivalent to $\mathcal{I}$ being $\omega$-maximal.
+Theorem 5.11.
+Let $\mathcal{I}$ be a proper ideal on $\mathbb{N}$. Then, the following statements are equivalent:
+  1. (1)
+    $\mathcal{I}$ is $\omega$-maximal.
+  2. (2)
+    There is a Banach lattice isomorphism from $\ell_{\infty}/c_{0,\mathcal{I}}$ into $\ell_{\infty}$.
+  3. (3)
+    There exists a bounded linear operator $Q\colon\ell_{\infty}\to\ell_{\infty}$ with $\ker Q=c_{0,\mathcal{I}}$ that verifies the next property: for every pair of subsets $A,B\subseteq\mathbb{N}$ and every $n\in\mathbb{N}$,
+    $Q(\chi_{A})_{n}\neq 0\quad\mbox{and}\quad Q(\chi_{B})_{n}\neq 0\;\Rightarrow\;Q(\chi_{A\cap B})_{n}\neq 0.$ (5.1)
+Proof.
+$(1)\Rightarrow(2)$ Notice that $\ell_{\infty}/c_{0,\mathcal{I}_{n}}$ is Banach lattice isomorphic to $\mathbb{R}$ for each $n\in\mathbb{N}$ via the map ${\bf x}+c_{0,\mathcal{I}_{n}}\mapsto\mathcal{I}_{n}-\lim{\bf x}$.
+Hence, $\ell_{\infty}((\ell_{\infty}/c_{0,\mathcal{I}_{n}})_{n\in\mathbb{N}})$ is Banach lattice isomorphic to $\ell_{\infty}$.
+Now the map $L\colon\ell_{\infty}/c_{0,\mathcal{I}}\to\ell_{\infty}((\ell_{\infty}/c_{0,\mathcal{I}_{n}})_{n\in\mathbb{N}})$, ${\bf x}+c_{0,\mathcal{I}}\mapsto({\bf x}+c_{0,\mathcal{I}_{n}})_{n\in\mathbb{N}}$ is a Banach lattice embedding.
+Thus, there is a Banach lattice isomorphism from $\ell_{\infty}/c_{0,\mathcal{I}}$ into $\ell_{\infty}$.
+$(2)\Rightarrow(3)$ Let $T\colon\ell_{\infty}/c_{0,\mathcal{I}}\to\ell_{\infty}$ be a Banach lattice embedding.
+Define $Q\colon\ell_{\infty}\to\ell_{\infty}$ by $Q({\bf x})=T({\bf x}+c_{0,\mathcal{I}})$ for each $\bf x$$\in\ell_{\infty}$.
+It is clear that $Q$ is linear and bounded.
+Since $T$ is injective, we have:
+$Q({\bf x})={\bf 0}\Longleftrightarrow T({\bf x}+c_{0,\mathcal{I}})={\bf 0}\Longleftrightarrow{\bf x}\in c_{0,\mathcal{I}}.$
+Thus, $\ker Q=c_{0,\mathcal{I}}$.
+We will verify (5.1).
+Let $n\in\mathbb{N}$ and $A,B\subseteq\mathbb{N}$ be such that $Q(\chi_{A})_{n}\neq 0$ and $Q(\chi_{B})_{n}\neq 0$.
+Observe that $Q(\chi_{A})=T(\chi_{A}+c_{0,\mathcal{I}})\geq{\bf 0}$ and $Q(\chi_{B})=T(\chi_{B}+c_{0,\mathcal{I}})\geq{\bf 0}$.
+Hence, $Q(\chi_{A})_{n}=T(\chi_{A}+c_{0,\mathcal{I}})_{n}>0$ and $Q(\chi_{B})_{n}=T(\chi_{B}+c_{0,\mathcal{I}})_{n}>0$.
+From the fact $(\chi_{A}+c_{0,\mathcal{I}})\wedge(\chi_{B}+c_{0,\mathcal{I}})=\chi_{A\cap B}+c_{0,\mathcal{I}}$, it follows that $T(\chi_{A\cap B}+c_{0,\mathcal{I}})=T(\chi_{A}+c_{0,\mathcal{I}})\wedge T(\chi_{A}+c_{0,\mathcal{I}})$ and
+$Q(\chi_{A\cap B})_{n}=T(\chi_{A\cap B}+c_{0,\mathcal{I}})_{n}=T(\chi_{A}+c_{0,\mathcal{I}})_{n}\wedge T(\chi_{A}+c_{0,\mathcal{I}})_{n}>0.$
+Thus, $Q(\chi_{A\cap B})_{n}\neq 0$.
+$(3)\Rightarrow(1)$ Let $Q\colon\ell_{\infty}\to c_{0,\mathcal{I}}$ be a bounded linear operator satisfying (5.1). For each $n\in\mathbb{N}$, consider the family
+$\mathcal{A}_{n}=\{A\subseteq\mathbb{N}:\;Q(\chi_{A})_{n}\neq 0\}$
+as in Lemma 5.10 with $\mathcal{J}=\mathcal{P}(\mathbb{N})$. Observe that $\mathcal{I}^{+}=\bigcup_{n\in\mathbb{N}}\mathcal{A}_{n}$ by Lemma 5.10. For each $n\in\mathbb{N}$, let
+$\mathcal{H}_{n}=\{B\subseteq\mathbb{N}:(\exists\,A\in\mathcal{A}_{n})(A\subseteq B)\}.$
+Notice that $\mathcal{H}_{n}\subseteq\mathcal{I}^{+}$ for all $n$.
+Indeed, suppose not and let $B\in\mathcal{H}_{n}$ with $B\not\in\mathcal{I}^{+}$.
+Let $A\in\mathcal{A}_{n}$ be such that $A\subseteq B$.
+Since $B\in\mathcal{I}$, $A\in\mathcal{I}$ and thus $Q(\chi_{A})={\bf 0}$, which is absurd.
+We claim that $\mathcal{I}_{n}\coloneqq 2^{\mathbb{N}}\setminus\mathcal{H}_{n}$ is an ideal.
+By construction, $\mathcal{I}_{n}$ is closed by taking subsets.
+Now, let $B,C\in\mathcal{I}_{n}$ and suppose that $B\cup C\not\in\mathcal{I}_{n}$, that is, $B\cup C\in\mathcal{H}_{n}$.
+Let $A\in\mathcal{A}_{n}$ be such that $A\subseteq B\cup C.$ Since $B\cap C\not\in\mathcal{A}_{n}$, by (5) of Lemma 5.10, we have either $A\cap B\in\mathcal{A}_{n}$ or $A\cap C\in\mathcal{A}_{n}$.
+It follows that either $B\in\mathcal{H}_{n}$ or $C\in\mathcal{H}_{n}$, a contradiction.
+So, $\mathcal{I}_{n}$ is an ideal.
+Now as $\mathcal{I}$ is proper, $\mathbb{N}\not\in\mathcal{I}$.
+Thus, $S=\{n\in\mathbb{N}\,\colon\,Q(\chi_{\mathbb{N}})_{n}\neq 0\}$ is non-empty.
+Observe that $\mathbb{N}\in\mathcal{A}_{n}$ if and only if $n\in S$.
+Therefore, $\mathcal{I}_{n}\neq\mathcal{P}(\mathbb{N})$ if and only if $n\in S$.
+By (4) of Lemma 5.10, $\mathcal{I}=\bigcap_{n}\mathcal{I}_{n}=\bigcap_{n\in S}\mathcal{I}_{n}$.
+It remains to show that each $\mathcal{I}_{n}$ is maximal for each $n\in S$.
+If not, let $B\subseteq\mathbb{N}$ be such that $B\not\in\mathcal{I}_{n}$ and $\mathbb{N}\setminus B\not\in\mathcal{I}_{n}$.
+Then there are $C_{0},C_{1}\in\mathcal{A}_{n}$ such that $C_{0}\subseteq B$ and $C_{1}\subseteq\mathbb{N}\setminus B$.
+So, $Q(\chi_{C_{0}\cap C_{1}})={\bf 0}$.
+On the other hand, by (5.1) we have $Q(\chi_{C_{0}\cap C_{1}})_{n}\neq 0$, an absurd.
+This completes the proof.
+∎
+Remark 5.12.
+The Property (5.1) cannot be omitted.
+Indeed, by Corollary 5.3, the existence of a bounded linear operator $Q$ from $\ell_{\infty}$ to $\ell_{\infty}$ with $\ker Q=c_{0,\mathcal{I}}$ implies that $\mathcal{I}$ is complemented.
+However, [13, Example 2] shows a complemented ideal which is not $\omega$-maximal.
+Observe that if $\mathcal{I}$ is a strongly $\omega$-maximal ideal, then $\mathcal{I}$ is complemented by Theorem 5.7.
+Our next result provides an alternative proof by constructing an explicit projection.
+To this end, we first establish an auxiliary result.
 Recall that if ${\bf x}=(x_{n})\in\ell_{\infty}$, its support is defined as $\mathrm{supp}({\bf x})\coloneqq\{n\in\mathbb{N}\,\colon\,x_{n}\neq 0\}$.
@@
 Then, the sum $\sum_{n\in F}{\bf x}_{n}$ is the sequence in $\ell_{\infty}$ whose $m$-th coordinate is ${\bf x}_{n}(m)$ whenever $m\in\mathrm{supp}({\bf x}_{n})$ for some $n\in F$, and 0 otherwise.
+Observe that the previous sum is well defined since the supports are pairwise disjoint.
 Lemma 5.13.
@@
 ∎
+Next result presents an explicit projection in the case of strongly $\omega$-maximal ideals.
 Theorem 5.14.
@@
   5. (5)
-    The family =$\mathcal{B}=\{B\subseteq\mathbb{N}\,\colon\,T(B)=B\}$ is closed under arbitrary intersections.
+    The family $\mathcal{B}=\{B\subseteq\mathbb{N}\,\colon\,T(B)=B\}$ is closed under arbitrary intersections.
 Proof.
@@
 ${\bf x}$ $\mapsto{\bf x}-\sum_{m\in\mathbb{N}}(\mathcal{I}_{m}^{*}-\lim{\bf x})\chi_{A_{m}}.$
-Now define $Q\colon\ell_{\infty}\to\ell_{\infty}$ by $Q=Id-P$, that is,
+Now define $Q\colon\ell_{\infty}\to\ell_{\infty}$ by Q=IdPQ=Id-P, that is,
 $Q({\bf x})=\sum_{m\in\mathbb{N}}(\mathcal{I}_{m}^{*}-\lim{\bf x})\chi_{A_{m}},\quad{\bf x}\in\ell_{\infty}.$ (5.2)
@@
 ∎
-Remark 5.10.
-Property (4) of Theorem 5.8 can be restated as follows: For any pair $A,B$ of subsets of $\mathbb{N}$ and $n\in\mathbb{N}$,
-$Q(\chi_{A})_{n}\neq 0\quad\mbox{and}\quad Q(\chi_{B})_{n}\neq 0\;\Rightarrow\;Q(\chi_{A\cap B})_{n}\neq 0.$ (5.2)
-We will show that, under some extra condition, the complementation of $\mathcal{I}$ implies that $\mathcal{I}$ is $\omega$-maximal.
-For that end, we need a key lemma whose argument is motivated by the proof of the non-complementation of $c_{0}$ in $\ell_{\infty}$ due to Whitley (see [7, Theorem 5.6] or [1, Theorem 2.5.4]).
-Indeed, we will show a more general fact: $c_{0}$ is not complemented in $c_{0,\mathcal{J}}$ for any ideal $\mathcal{J}\supsetneq{\sf Fin}$.
-A similar result appears in [18, Corollary 1.5] about the space $c$ of convergent sequences.
-Lemma 5.11.
-Let $\mathcal{I}\subsetneq\mathcal{J}$ be ideals on $\mathbb{N}$ such that $c_{0,\mathcal{I}}$ is complemented in $c_{0,\mathcal{J}}$.
-Then, there exists a collection $\{\mathcal{A}_{n,k}\colon\,n,k\in\mathbb{N}\}$ of subsets of $2^{\mathbb{N}}$ satisfying:
-  1. (1)
-    $\mathcal{A}_{n,k}\subseteq\mathcal{J}\cap\mathcal{I}^{+}$ for each $n,k\in\mathbb{N}$.
-  2. (2)
-    $\mathcal{A}_{n,k}\subseteq\mathcal{A}_{n,k+1}$ for all $n,k\in\mathbb{N}$.
-  3. (3)
-    There is $M>0$ such that $\mathfrak{ad}_{\mathcal{I}}(\mathcal{A}_{n,k})\leq kM$ for all $n,k\in\mathbb{N}$.
-  4. (4)
-    For each $n\in\mathbb{N}$, let $\mathcal{A}_{n}=\bigcup_{k\in\mathbb{N}}\mathcal{A}_{n,k}$. Then $\mathcal{J}\cap\mathcal{I}^{+}=\bigcup_{n\in\mathbb{N}}\mathcal{A}_{n}$.
-  5. (5)
-    If $A\in\mathcal{A}_{n,k}$ and $A=B\cup C$ with $B\cap C\not\in\mathcal{A}_{n}$, then either $B\in\mathcal{A}_{n,2k}$ or $C\in\mathcal{A}_{n,2k}$.
-In particular, if $A\in\mathcal{A}_{n}$ and $A=B\cup C$ with $B\cap C\not\in\mathcal{A}_{n}$, then either $B\in\mathcal{A}_{n}$ or $C\in\mathcal{A}_{n}$.
-Proof.
-Let $P\colon c_{0,\mathcal{J}}\to c_{0,\mathcal{I}}$ be a projection onto $c_{0,\mathcal{I}}$ and $Q=Id-P$. For each $n,k\in\mathbb{N}$, define
-$\mathcal{A}_{n,k}=\{A\in\mathcal{J}\,\colon\,|Q(\chi_{A})_{n}|\geq 1/k\}.$
-Since =$\mathcal{I}=\{A\in\mathcal{J}\,\colon\,Q(\chi_{A})={\bf 0}\}$, we have
-$\mathcal{J}\cap\mathcal{I}^{+}$ =$=\{A\in\mathcal{J}\,\colon\,Q(\chi_{A})\neq{\bf 0}\}$
-$=\bigcup_{k,n\in\mathbb{N}}\{A\in\mathcal{J}\,\colon\,|Q(\chi_{A})_{n}|\geq 1/k\}$
-$=\bigcup_{n\in\mathbb{N}}\mathcal{A}_{n}.$
-Then (1) and (4) hold.
-(2) is obvious.
-To see (3), let $m\in\mathbb{N}$ be given and $A_{1},\ldots,A_{m}\in\mathcal{A}_{n,k}$ be such that $A_{i}\cap A_{j}\in\mathcal{I}$ for $i\neq j$.
-If $F_{1}=A_{1}$ and $F_{j}=A_{j}\setminus(A_{1}\cup\cdots\cup A_{j-1})$ for $2\leq j\leq m$, then $Q(\chi_{A_{j}})=Q(\chi_{F_{j}})$ for all $1\leq j\leq m$.
-For $1\leq j\leq m$, let $a_{j}\in\mathbb{K}$ be such that $a_{j}Q(\chi_{F_{j}})_{n}=|Q(\chi_{F_{j}})_{n}|$.
-Thus,
-$\frac{m}{k}\leq\sum_{j=1}^{m}|Q(\chi_{A_{j}})_{n}|$ $=\sum_{j=1}^{m}|Q(\chi_{F_{j}})_{n}|$
-$=\sum_{j=1}^{m}a_{j}Q(\chi_{F_{j}})_{n}$
-$\leq\left\|\sum_{j=1}^{m}a_{j}Q(\chi_{F_{j}})\right\|$
-$\leq\|Q\|\left\|\sum_{j=1}^{m}a_{j}\chi_{F_{j}}\right\|=\|Q\|.$
-Hence, $m\leq k\|Q\|$.
-Finally, we show (5).
-Let $A\in\mathcal{A}_{n,k}$ and $A=B\cup C$ with $B\cap C\not\in\mathcal{A}_{n}$.
-Then $|Q(\chi_{B\cap C})_{n}|=0$.
-Assume that $|Q(\chi_{B})_{n}|<1/2k$ and $|Q(\chi_{C})_{n}|<1/2k$.
-Note that $Q(\chi_{B})_{n}=Q(\chi_{B\setminus C})_{n}$ and $Q(\chi_{C})_{n}=Q(\chi_{C\setminus B})_{n}$.
-Thus
-$|Q(\chi_{A})_{n}|=|Q(\chi_{A\setminus B\cap C})_{n}|=|Q(\chi_{B})_{n}+Q(\chi_{C})_{n}|\leq 1/k,$
-which is absurd. ∎
-Theorem 5.12.
-Let $\mathcal{I}\subsetneq\mathcal{J}$ be proper ideals on $\mathbb{N}$ such that $c_{0,\mathcal{I}}$ is complemented in $c_{0,\mathcal{J}}$.
-Then $\mathfrak{ad}_{\mathcal{I}}(\mathcal{J}\cap\mathcal{I}^{+})\leq\aleph_{0}$.
-Proof.
-For $k,n\in\mathbb{N}$, let $\mathcal{A}_{n,k}$ be as in Lemma 5.11.
-Then, we have $\mathcal{J}\cap\mathcal{I}^{+}=\bigcup_{n,k}\mathcal{A}_{n,k}$ and $\mathfrak{ad}_{\mathcal{I}}(\mathcal{A}_{n,k})<\aleph_{0}$ for each $n,k$.
-Suppose $\mathcal{B}\subseteq\mathcal{J}\cap\mathcal{I}^{+}$ is an $\mathcal{I}$-AD family.
-Then, for every $n,k\in\mathbb{N}$, $\mathcal{B}\cap\mathcal{A}_{n,k}$ is finite.
-Thus, $\mathcal{B}$ is countable, and consequently $\mathfrak{ad}_{\mathcal{I}}(\mathcal{J}\cap\mathcal{I}^{+})\leq\aleph_{0}$.
-∎
-The next observation follows immediately from Lemma 5.1.
-Corollary 5.13.
-Let $\mathcal{I}\subsetneq\mathcal{J}$ be proper ideals on $\mathbb{N}$ such that $\mathcal{I}\restriction A$ is meager as a subset of $2^{A}$ for some infinite set $A\in\mathcal{J}$.
-Then $\mathfrak{ad}_{\mathcal{I}}(\mathcal{J}\cap\mathcal{I}^{+})>\aleph_{0}$.
-An ideal $\mathcal{I}$ is everywhere meager [8], if $\mathcal{I}\restriction A$ is meager in $2^{A}$ for all $A\in\mathcal{I}^{+}$, which in turns is equivalent to requiring that $\mathcal{I}\restriction A$ is Baire measurable in $2^{A}$ for all $A\in\mathcal{I}^{+}$.
-Every analytic ideal is everywhere meager.
-On the opposite side, a complemented ideal $\mathcal{I}$ is nowhere meager, since $\mathcal{I}\restriction A$ is not meager in $2^{A}$ for all $A\in\mathcal{I}^{+}$ by Proposition 5.6.
-From Theorem 5.12 we get the following.
-Corollary 5.14.
-Let $\mathcal{I}\subsetneq\mathcal{J}$ be proper ideals on $\mathbb{N}$ such that $\mathcal{I}$ is everywhere meager.
-Then $c_{0,\mathcal{I}}$ is not complemented in $c_{0,\mathcal{J}}$.
-In particular, $c_{0}$ is not complemented in $c_{0,\mathcal{J}}$ for any $\mathcal{J}\supsetneq{\sf Fin}$.
-Remark 5.15.
-Concerning the above results, Theorem 4.6 gives examples of proper ideals $\mathcal{I}$ and $\mathcal{J}$ with ${\sf Fin}\subsetneq\mathcal{I}\subsetneq\mathcal{J}$ such that $c_{0,\mathcal{I}}$ is complemented in $c_{0,\mathcal{J}}$.
-Indeed, if $\mathcal{J}$ is a proper ideal and $\mathcal{I}_{1},\ldots,\mathcal{I}_{k}$ are maximal ideals such that $\mathcal{J}\not\subseteq\mathcal{I}_{j}$ for all $1\leq j\leq k$, then $\mathcal{I}\coloneqq\mathcal{J}\cap(\bigcap_{1\leq j\leq k}\mathcal{I}_{j})$ is $k$-maximal in $\mathcal{J}$.
-Thus, $c_{0,\mathcal{I}}$ is complemented in $c_{0,\mathcal{J}}$ by Theorem 4.6.
-Notice that in this case $\dim(c_{0,\mathcal{J}}/c_{0,\mathcal{I}})=k$.
-We do not know if there exist examples with $\dim(c_{0,\mathcal{J}}/c_{0,\mathcal{I}})=\infty$.
-In spite of the previous results, it is still possible for $c_{0,\mathcal{I}}$ to contain complemented subspaces that are isomorphic or isometric to $c_{0}$.
-In particular, we proved in [23] that $\ell_{\infty}(c_{0})$ is isometric to $c_{0,{\sf Fin}^{\omega}}$, and since $\ell_{\infty}(c_{0})$ clearly contains a complemented copy of $c_{0}$, the same holds for $c_{0,{\sf Fin}^{\omega}}$.
-This contrasts with the situation in $\ell_{\infty}$, where every subspace isomorphic or isometric to $c_{0}$ is necessarily not complemented.
-This leads to the following result.
-Proposition 5.16.
-Let $\mathcal{I}$ be an ideal on $\mathbb{N}$. If $c_{0,\mathcal{I}}$ has a complemented copy of $c_{0}$, then $\mathcal{I}$ is not complemented.
-Proof.
-Suppose that $\mathcal{I}$ is complemented, that is, $\ell_{\infty}=c_{0,\mathcal{I}}\oplus W$.
-Since $c_{0,\mathcal{I}}$ has a complemented copy of $c_{0}$, there is a subspace $E$ isomorphic to $c_{0}$ such that $c_{0,\mathcal{I}}=E\oplus Z$.
-Thus, $E$ is complemented in $\ell_{\infty}$, which is impossible by Lindenstrauss’s theorem.
-∎
-However, it is unclear whether $c_{0,\mathcal{I}}$ contains a complemented copy of $c_{0}$ for any meager ideal $\mathcal{I}$.
-The next corollary is due to Kania [14, Theorem A].
-Corollary 5.17.
-Let $\mathcal{I}$ be an ideal on $\mathbb{N}$. If $\mathcal{I}$ is complemented, then $\mathfrak{ad}(\mathcal{I}^{+})\leq\aleph_{0}$.
-Now we prove, with an extra assumption, that the complementation of $c_{0,\mathcal{I}}$ implies that $\mathcal{I}$ is $\omega$-maximal.
-Theorem 5.18.
-Let $\mathcal{I}$ be a proper ideal on $\mathbb{N}$.
-Suppose $c_{0,\mathcal{I}}$ is complemented in $\ell_{\infty}$ by a projection map $P\colon\ell_{\infty}\to c_{0,\mathcal{I}}$ such that $Q:=Id-P$ satisfies (5.2).
-Then $\mathcal{I}$ is $\omega$-maximal.
-Proof.
-Suppose $P\colon\ell_{\infty}\to c_{0,\mathcal{I}}$ is a projection map as in the hypothesis. For each $n\in\mathbb{N}$ consider the family
-$\mathcal{A}_{n}=\{A\subseteq\mathbb{N}:\;Q(\chi_{A})_{n}\neq 0\}$
-as in Lemma 5.11 with $\mathcal{J}=\mathcal{P}(\mathbb{N})$. Observe that $\mathcal{I}^{+}=\bigcup_{n\in\mathbb{N}}\mathcal{A}_{n}$ by Lemma 5.11. For each $n\in\mathbb{N}$, let
-$\mathcal{H}_{n}=\{B\subseteq\mathbb{N}:(\exists\,A\in\mathcal{A}_{n})(A\subseteq B)\}.$
-Notice that $\mathcal{H}_{n}\subseteq\mathcal{I}^{+}$ for all $n$.
-Indeed, suppose not and let $B\in\mathcal{H}_{n}$ with $B\not\in\mathcal{I}^{+}$.
-Let $A\in\mathcal{A}_{n}$ be such that $A\subseteq B$.
-Since $B\in\mathcal{I}$, $A\in\mathcal{I}$ and thus $Q(\chi_{A})={\bf 0}$, which is absurd.
-We claim that $\mathcal{I}_{n}\coloneqq 2^{\mathbb{N}}\setminus\mathcal{H}_{n}$ is an ideal.
-By construction, $\mathcal{I}_{n}$ is closed by taking subsets.
-Now, let $B,C\in\mathcal{I}_{n}$ and suppose that $B\cup C\not\in\mathcal{I}_{n}$, that is, $B\cup C\in\mathcal{H}_{n}$.
-Let $A\in\mathcal{A}_{n}$ be such that $A\subseteq B\cup C.$ Since $B\cap C\not\in\mathcal{A}_{n}$, by (5) of Lemma 5.11, we have either $A\cap B\in\mathcal{A}_{n}$ or $A\cap C\in\mathcal{A}_{n}$.
-It follows that either $B\in\mathcal{H}_{n}$ or $C\in\mathcal{H}_{n}$, a contradiction.
-So, $\mathcal{I}_{n}$ is an ideal.
-Now as $\mathcal{I}$ is proper, $\mathbb{N}\not\in\mathcal{I}$.
-Thus, $S=\{n\in\mathbb{N}\,\colon\,Q(\chi_{\mathbb{N}})_{n}\neq 0\}$ is non-empty.
-Observe that $\mathbb{N}\in\mathcal{A}_{n}$ if and only if $n\in S$.
-Therefore, $\mathcal{I}_{n}\neq\mathcal{P}(\mathbb{N})$ if and only if $n\in S$.
-By (4) of Lemma 5.11, $\mathcal{I}=\bigcap_{n}\mathcal{I}_{n}=\bigcap_{n\in S}\mathcal{I}_{n}$.
-It remains to show that each $\mathcal{I}_{n}$ is maximal for each $n\in S$.
-If not, let $B\subseteq\mathbb{N}$ be such that $B\not\in\mathcal{I}_{n}$ and $\mathbb{N}\setminus B\not\in\mathcal{I}_{n}$.
-Then there are $C_{0},C_{1}\in\mathcal{A}_{n}$ such that $C_{0}\subseteq B$ and $C_{1}\subseteq\mathbb{N}\setminus B$.
-So, $Q(\chi_{C_{0}\cap C_{1}})={\bf 0}$.
-On the other hand, by (5.2) we have $Q(\chi_{C_{0}\cap C_{1}})_{n}\neq 0$, an absurd.
-∎
-Remark 5.19.
-It is not difficult to verify that $\mathcal{I}_{n}$ is maximal if and only if Property (5.2) holds for that $n$. We do not know whether the assumption of Property (5.2) can be removed.
-Theorem 5.18 raises the following question: Is there an $\omega$-maximal ideal that is not complemented?
-In other words, is being complemented equivalent to being $\omega$-maximal?
-A natural candidate for a negative answer to this question is the $\omega$-maximal ideal constructed in Example 3.8.
 Our next result is the converse of Theorem 5.14.
@@
   • If $A,B\subseteq\mathbb{N}$, then $T(A\cap B)=T(A)\cap T(B)$.
-  • The family =$\mathcal{B}=\{B\subseteq\mathbb{N}\,\colon\,T(B)=B\}$ is closed under arbitrary intersections.
+  • The family $\mathcal{B}=\{B\subseteq\mathbb{N}\,\colon\,T(B)=B\}$ is closed under arbitrary intersections.
 Then $\mathcal{I}$ is strongly $\omega$-maximal. Moreover, there is a partition of $\mathbb{N}$ such that $Q$ has the same form as given in (5.2).
@@
 Thus, $Q({\bf x})=\sum_{m\in F}(\mathcal{I}_{m}^{*}-\lim{\bf x})\chi_{V_{m}}$ when ${\bf x}=\sum_{i=1}^{k}c_{i}\chi_{A_{i}}$.
-The general case follows since the linear span of the set $\{\chi_{A}\,\colon\,A\subseteq\mathbb{N}\}$ is dense in $\ell_{\infty}$ (see [23, Proposition 2.5]).
+The general case follows since the linear span of the set $\{\chi_{A}\,\colon\,A\subseteq\mathbb{N}\}$ is dense in $\ell_{\infty}$ (see [25, Proposition 2.5]).
 Therefore, $Q$ has the same form as given in (5.2).
 ∎
+Now, in the same spirit of Theorems 5.2 and 5.11, we give a characterization of strongly $\omega$-maximal ideals in terms of the quotient $\ell_{\infty}/c_{0,\mathcal{I}}$.
+Theorem 5.19.
+Let $\mathcal{I}$ be an ideal on $\mathbb{N}$. The following statements are equivalent:
+  1. (1)
+    $K_{\mathcal{I}}$ is homeomorphic to $\beta\mathbb{N}$;
+  2. (2)
+    $\ell_{\infty}/c_{0,\mathcal{I}}$ is isometric to $\ell_{\infty}$;
+  3. (3)
+    $\ell_{\infty}/c_{0,\mathcal{I}}$ is Banach lattice isomorphic to $\ell_{\infty}$;
+  4. (4)
+    $\mathcal{I}$ is strongly $\omega$-maximal.
+  5. (5)
+    $c_{0,\mathcal{I}}$ is strongly complemented.
+Proof.
+$(1)\Longleftrightarrow(2)$ It is classical result that $C(\beta\mathbb{N})$ is isometric to $\ell_{\infty}$ and $C(K_{\mathcal{I}})$ is isometric to $\ell_{\infty}/c_{0,\mathcal{I}}$ by Theorem 5.1.
+Thus, the claim follows from the classical Banach-Stone theorem [1, Theorem 4.1.5].
+$(1)\Longleftrightarrow(3)$ is shown analogously using a well known result of Kaplansky [17].
+$(4)\Longleftrightarrow(5)$ is proved in Theorems 5.14 and 5.16.
+$(4)\Rightarrow(3)$.
+Suppose that $\mathcal{I}$ is strongly $\omega$-maximal.
+By Proposition 3.1 and Lemma 2.3, there is a partition $\{A_{n}\,\colon\,n\in\mathbb{N}\}$ such that $A_{n}\in\mathcal{I}_{n}^{*}$ for every $n\in\mathbb{N}$ and $A\in\mathcal{I}$ if and only if $A\cap A_{n}\in\mathcal{I}_{n}$ for each $n\in\mathbb{N}$.
+Claim 5.20.
+The map
+$\Psi\colon\ell_{\infty}$ $\to\ell_{\infty}((\ell_{\infty}/c_{0,\mathcal{I}_{m}})_{m\in\mathbb{N}})$
+${\bf x}=(x_{n})$ $\mapsto((x_{n}\chi_{A_{m}}(n))+c_{0,\mathcal{I}_{m}})$
+is an onto Banach lattice homomorphism and $\ker\Psi=c_{0,\mathcal{I}}$.
+In fact, from the definition, it is clear that $\Psi$ is linear and preserves the lattice operations.
+To check continuity, we will use the following formula proved in [25, Lemma 5.7].
+For ${\bf x}=(x_{n})\in\ell_{\infty}$ we have
+$\|\Psi({\bf x})\|=\sup_{m\in\mathbb{N}}\|(x_{n}\chi_{A_{m}}(n))+c_{0,\mathcal{I}_{m}}\|=\sup_{m\in\mathbb{N}}\bigg(\mathcal{I}_{m}-\lim|x_{n}\chi_{A_{m}}(n)|\bigg)\leq\|{\bf x}\|.$
+We will verify that $\Psi$ is onto.
+Let $({\bf y}^{m}+c_{0,\mathcal{I}_{m}})\in\ell_{\infty}((\ell_{\infty}/c_{0,\mathcal{I}_{m}})_{m\in\mathbb{N}})$ be given.
+For each $m\in\mathbb{N}$, let $a_{m}=\mathcal{I}_{m}-\lim{\bf y}^{m}$, which exists because $\mathcal{I}_{m}$ is maximal.
+Then, for every $m\in\mathbb{N}$,
+${\bf y}^{m}+c_{0,\mathcal{I}_{m}}=a_{m}{\bf 1}+c_{0,\mathcal{I}_{m}}=a_{m}\chi_{A_{m}}+c_{0,\mathcal{I}_{m}}.$
+Define ${\bf x}=(x_{n})\in\ell_{\infty}$ by $x_{n}=a_{m}$ whenever $n\in A_{m}$. Hence,
+$\Psi({\bf x})=((x_{n}\chi_{A_{m}}(n))+c_{0,\mathcal{I}_{m}})=((a_{m}\chi_{A_{m}}(n))+c_{0,\mathcal{I}_{m}})=({\bf y}^{m}+c_{0,\mathcal{I}_{m}}).$
+and therefore $\Psi$ is surjective.
+Next, we prove that $\ker\Psi=c_{0,\mathcal{I}}$.
+Suppose that ${\bf x}=(x_{n})\in\ell_{\infty}$ satisfies that $\Psi({\bf x})={\bf 0}$.
+Then $((x_{n}\chi_{A_{m}}(n))\in c_{0,\mathcal{I}_{m}})$ for each $m\in\mathbb{N}$.
+Let $\varepsilon>0$ be given.
+By definition, $A(((x_{n}\chi_{A_{m}}(n)),\varepsilon)=A({\bf x},\varepsilon)\cap A_{m}\in\mathcal{I}_{m}$ for every $m\in\mathbb{N}$.
+By Lemma 2.3, it follows that $A({\bf x},\varepsilon)\in\mathcal{I}$.
+Since $\varepsilon>0$ was arbitrary, we conclude that ${\bf x}\in c_{0,\mathcal{I}}$.
+Finally, the map $\Psi$ induces a Banach lattice isomorphism from $\ell_{\infty}/c_{0,\mathcal{I}}$ onto $\ell_{\infty}((\ell_{\infty}/c_{0,\mathcal{I}_{m}})_{m\in\mathbb{N}})$.
+Since $\dim\ell_{\infty}/c_{0,\mathcal{I}_{m}}=1$ for each $m\in\mathbb{N}$, we have that $\ell_{\infty}/c_{0,\mathcal{I}_{m}}$ is Banach lattice isomorphic to $\mathbb{R}$.
+By Remark 5.4, $\ell_{\infty}((\ell_{\infty}/c_{0,\mathcal{I}_{m}})_{m\in\mathbb{N}})$ is Banach lattice isomorphic to $\ell_{\infty}$.
+Consequently, $\ell_{\infty}/c_{0,\mathcal{I}}$ is Banach lattice isomorphic to $\ell_{\infty}$.
+$(3)\Longrightarrow(4)$ Suppose that $\ell_{\infty}/c_{0,\mathcal{I}}$ is Banach lattice isomorphic to $\ell_{\infty}$ and let $T\colon\ell_{\infty}/c_{0,\mathcal{I}}\to\ell_{\infty}$ be a Banach lattice isomorphism.
+For each $n\in\mathbb{N}$, let ${\bf x}_{n}\in\ell_{\infty}$ be such that $T({\bf x}_{n}+c_{0,\mathcal{I}})=e_{n}$.
+Since ${\bf x}_{n}\not\in c_{0,\mathcal{I}}$, there are $\varepsilon_{n}>0$ and $A_{n}\not\in\mathcal{I}$ satisfying $\varepsilon_{n}\chi_{A_{n}}\leq|{\bf x}_{n}|$.
+Thus, $T(\varepsilon_{n}\chi_{A_{n}}+c_{0,\mathcal{I}})\leq T({\bf x}_{n}+c_{0,\mathcal{I}})=e_{n}$.
+Hence, $T(\chi_{A_{n}}+c_{0,\mathcal{I}})=a_{n}e_{n}$ for some $a_{n}>0$.
+Claim 5.21.
+If $m,n\in\mathbb{N}$ and $m\neq n$, then $A_{n}\cap A_{m}\in\mathcal{I}$.
+Indeed, we have $\chi_{A_{m}\cap A_{n}}+c_{0,\mathcal{I}}=(\chi_{A_{m}}+c_{0,\mathcal{I}})\wedge(\chi_{A_{n}}+c_{0,\mathcal{I}})$. Hence,
+$T(\chi_{A_{m}\cap A_{n}}+c_{0,\mathcal{I}})=T(\chi_{A_{m}}+c_{0,\mathcal{I}})\wedge T(\chi_{A_{n}}+c_{0,\mathcal{I}})=a_{m}e_{m}\wedge a_{n}e_{n}={\bf 0}.$
+Claim 5.22.
+For each $n\in\mathbb{N}$, $\mathcal{I}\restriction A_{n}$ is maximal.
+Fix $n\in\mathbb{N}$.
+Let $B\subseteq A_{n}$ be such that $B\not\in\mathcal{I}$.
+Then, $T(\chi_{B}+c_{0,\mathcal{I}})\leq T(\chi_{A_{n}}+c_{0,\mathcal{I}})=a_{n}e_{n}$.
+Thus, $T(\chi_{B}+c_{0,\mathcal{I}})=b_{n}e_{n}$ for some $0<b_{n}\leq a_{n}$.
+Consequently, $T((\chi_{A_{n}}-\frac{a_{n}}{b_{n}}\chi_{B})+c_{0,\mathcal{I}})={\bf 0}$.
+It follows that $A_{n}\setminus B\in\mathcal{I}$ because of $A_{n}\setminus B\subset A({\bf z},1/2)$ where ${\bf z}=\chi_{A_{n}}-\frac{a_{n}}{b_{n}}\chi_{B}\in c_{0,\mathcal{I}}$.
+Claim 5.23.
+$B\in\mathcal{I}$ if, and only if, $B\cap A_{n}\in\mathcal{I}$ for each $n\in\mathbb{N}$.
+Suppose that $B\cap A_{n}\in\mathcal{I}$ for each $n\in\mathbb{N}$ and let $T(\chi_{B}+c_{0,\mathcal{I}})={\bf y}=(y_{n})$. Fix $m\in\mathbb{N}$. Notice that ${\bf y}\geq 0$, and
+${\bf 0}=T(\chi_{B\cap A_{m}}+c_{0,\mathcal{I}})=T(\chi_{B}+c_{0,\mathcal{I}})\wedge T(\chi_{A_{m}}+c_{0,\mathcal{I}})={\bf y}\wedge(a_{m}e_{m})=\min\{y_{m},a_{m}\}e_{m}.$
+Thus, $y_{m}=0$. Since $m\in\mathbb{N}$ was arbitrary, we conclude that ${\bf y}={\bf 0}$. So, $B\in\mathcal{I}$.
+Now, following the argument used in the proof of Lemma 4.5, we may assume that $A_{m}\cap A_{n}=\emptyset$ for every $m,n\in\mathbb{N}$ with $m\neq n$.
+Define $\mathcal{I}_{n}=\mathcal{I}\restriction A_{n}\sqcup\mathcal{P}(A_{n}^{c})$ for each $n\in\mathbb{N}$.
+By (1) of Lemma 2.2 and Claim 5.22, $\mathcal{I}_{n}$ is maximal for all $n\in\mathbb{N}$.
+It follows from Claim 5.23 that $\mathcal{I}=\bigcap_{n\in\mathbb{N}}\mathcal{I}_{n}$.
+Therefore, by Proposition 3.1, $\mathcal{I}$ is strongly $\omega$-maximal.
+∎
+6. Relative complementation and AD-families
+In this section, we study when $c_{0,\mathcal{I}}$ is complemented in $c_{0,\mathcal{J}}$ when $\mathcal{I}\subseteq\mathcal{J}$.
+Also, we discuss the existence of $\mathcal{I}$-AD families and their connection with the complementation of $\mathcal{I}$.
+Recall that a $\mathcal{A}\subseteq\mathcal{I}^{+}$ is $\mathcal{I}$-AD if $X\cap Y\in\mathcal{I}$ for all $X,Y\in\mathcal{A}$. We let
+$\mathfrak{ad}_{\mathcal{I}}(\mathcal{A})=\max\{|\mathcal{B}|\,\colon\,\mbox{$\mathcal{B}\subseteq\mathcal{A}$ is an $\mathcal{I}$-AD family}\}.$
+We will write just $\mathfrak{ad}(\mathcal{A})$ when it is clear from the context which ideal $\mathcal{I}$ is used.
+We begin by stating a lemma, which is a well-known consequence of Theorem 2.1 (for a proof, see, for instance, [20]).
+Lemma 6.1.
+Let $\mathcal{I}$ be a meager ideal on $\mathbb{N}$. Then $\mathfrak{ad}(\mathcal{\mathcal{I}}^{+})=2^{\aleph_{0}}$.
+On the other hand, $\mathfrak{ad}(\mathcal{\mathcal{I}}^{+})\leq\aleph_{0}$ for every $\omega$-maximal ideal $\mathcal{I}$, this was implicitly shown by Plewik [24] and, for the sake of completion, we include a direct proof.
+Let $\mathcal{I}=\bigcap_{n}\mathcal{I}_{n}$ be an $\omega$-maximal ideal and suppose $\mathcal{B}\subseteq\mathcal{I}^{+}$ is an uncountable $\mathcal{I}$-AD family.
+Then, there is $n$ such that $|\mathcal{B}\cap\mathcal{I}_{n}^{+}|\geq 2$.
+If $A,B\in\mathcal{B}\cap\mathcal{I}_{n}^{+}$ are two different sets, then $A\cap B\in\mathcal{I}\subseteq\mathcal{I}_{n}$, which contradicts that $\mathcal{I}_{n}$ is maximal.
+Additionally, any $\omega$-maximal ideal is not measurable with respect to the usual product measure on $2^{\mathbb{N}}$ (see [16, section 4.1]).
+The proof presented above shows that:
+Corollary 6.2.
+(Kania [15, Theorem A]) Let $\mathcal{I}$ be an ideal on $\mathbb{N}$. If $\mathcal{I}$ is complemented, then $\mathfrak{ad}(\mathcal{I}^{+})\leq\aleph_{0}$.
+The following theorem extends Kania’s result to ideals $\mathcal{I}$ and $\mathcal{J}$ on $\mathbb{N}$ such that $\mathcal{I}\subsetneq\mathcal{J}$ and $c_{0,\mathcal{I}}$ complemented in $c_{0,\mathcal{J}}$.
+Theorem 6.3.
+Let $\mathcal{I}\subsetneq\mathcal{J}$ be proper ideals on $\mathbb{N}$.
+Suppose that $c_{0,\mathcal{I}}$ is kernel of an operator from $c_{0,\mathcal{J}}$ to $c_{0,\mathcal{J}}$.
+Then, $\mathfrak{ad}_{\mathcal{I}}(\mathcal{J}\cap\mathcal{I}^{+})\leq\aleph_{0}$.
+In particular, if $c_{0,\mathcal{I}}$ is complemented in $c_{0,\mathcal{J}}$, then $\mathfrak{ad}_{\mathcal{I}}(\mathcal{J}\cap\mathcal{I}^{+})\leq\aleph_{0}$.
+Proof.
+Let $Q\colon c_{0,\mathcal{J}}\to c_{0,\mathcal{J}}$ be a bounded linear operator such that $\ker Q=c_{0,\mathcal{I}}$.
+For $k,n\in\mathbb{N}$, let $\mathcal{A}_{n,k}$ be as in Lemma 5.10.
+Then, we have $\mathcal{J}\cap\mathcal{I}^{+}=\bigcup_{n,k}\mathcal{A}_{n,k}$ and $\mathfrak{ad}_{\mathcal{I}}(\mathcal{A}_{n,k})<\aleph_{0}$ for each $n,k$.
+Suppose $\mathcal{B}\subseteq\mathcal{J}\cap\mathcal{I}^{+}$ is an $\mathcal{I}$-AD family.
+Then, for every $n,k\in\mathbb{N}$, $\mathcal{B}\cap\mathcal{A}_{n,k}$ is finite.
+Thus, $\mathcal{B}$ is countable, and consequently $\mathfrak{ad}_{\mathcal{I}}(\mathcal{J}\cap\mathcal{I}^{+})\leq\aleph_{0}$.
+∎
+The next observation follows immediately from Lemma 6.1.
+Corollary 6.4.
+Let $\mathcal{I}\subsetneq\mathcal{J}$ be proper ideals on $\mathbb{N}$ such that $\mathcal{I}\restriction A$ is meager as a subset of $2^{A}$ for some infinite set $A\in\mathcal{J}$.
+Then, $\mathfrak{ad}_{\mathcal{I}}(\mathcal{J}\cap\mathcal{I}^{+})>\aleph_{0}$.
+An ideal $\mathcal{I}$ is everywhere meager [8], if $\mathcal{I}\restriction A$ is meager in $2^{A}$ for all $A\in\mathcal{I}^{+}$, which in turns is equivalent to requiring that $\mathcal{I}\restriction A$ is Baire measurable in $2^{A}$ for all $A\in\mathcal{I}^{+}$.
+Every analytic ideal is everywhere meager.
+On the opposite side, a complemented ideal $\mathcal{I}$ is nowhere meager, since $\mathcal{I}\restriction A$ is not meager in $2^{A}$ for all $A\in\mathcal{I}^{+}$ by Proposition 5.9.
+From Theorem 6.3 we get the following corollary. A similar result appears in [20, Corollary 1.5] about the space $c$ of convergent sequences.
+Corollary 6.5.
+Let $\mathcal{I}\subsetneq\mathcal{J}$ be proper ideals on $\mathbb{N}$ such that $\mathcal{I}$ is everywhere meager.
+Then $c_{0,\mathcal{I}}$ is not complemented in $c_{0,\mathcal{J}}$.
+In particular, $c_{0}$ is not complemented in $c_{0,\mathcal{J}}$ for any $\mathcal{J}\supsetneq{\sf Fin}$.
+Proposition 6.6.
+Let $\mathcal{I}$ and $\mathcal{J}$ be ideals on $\mathbb{N}$ with $\mathcal{I}\subsetneq\mathcal{J}$ and $k$ be a positive integer.
+If $\mathcal{J}$ is complemented and $\mathcal{I}$ is $k$-maximal in $\mathcal{J}$, then $\mathcal{I}$ is complemented.
+Proof.
+Since $\mathcal{I}$ is $k$-maximal in $\mathcal{J}$, it follows that $c_{0,\mathcal{I}}$ is complemented in $c_{0,\mathcal{J}}$ by Theorem 4.6.
+Additionally, $c_{0,\mathcal{J}}$ is complemented in $\ell_{\infty}$.
+Therefore, $c_{0,\mathcal{I}}$ is complemented in $\ell_{\infty}$.
+This completes the proof.
+∎
 Remark 6.7.
-We know that if $\mathcal{I}$ is an $\omega$-maximal ideal, then $\mathfrak{ad}(\mathcal{I}^{+})\leq\aleph_{0}$, as demonstrated earlier in the introduction to this section (see also [22]).
-The proof presented here shows also that if $\mathcal{I}$ is the intersection of a countable family of complemented ideals, then $\mathfrak{ad}(\mathcal{I}^{+})\leq\aleph_{0}$.
+Concerning the above results, Theorem 4.6 gives examples of proper ideals $\mathcal{I}$ and $\mathcal{J}$ with ${\sf Fin}\subsetneq\mathcal{I}\subsetneq\mathcal{J}$ such that $c_{0,\mathcal{I}}$ is complemented in $c_{0,\mathcal{J}}$.
+Indeed, if $\mathcal{J}$ is a proper ideal and $\mathcal{I}_{1},\ldots,\mathcal{I}_{k}$ are maximal ideals such that $\mathcal{J}\not\subseteq\mathcal{I}_{j}$ for all $1\leq j\leq k$, then $\mathcal{I}\coloneqq\mathcal{J}\cap(\bigcap_{1\leq j\leq k}\mathcal{I}_{j})$ is $k$-maximal in $\mathcal{J}$.
+Thus, $c_{0,\mathcal{I}}$ is complemented in $c_{0,\mathcal{J}}$ by Theorem 4.6.
+Notice that in this case $\dim(c_{0,\mathcal{J}}/c_{0,\mathcal{I}})=k$.
+We do not know if there exist examples with $\dim(c_{0,\mathcal{J}}/c_{0,\mathcal{I}})=\infty$.
+In spite of the previous results, it is still possible for $c_{0,\mathcal{I}}$ to contain complemented subspaces that are isomorphic or isometric to $c_{0}$.
+In particular, we proved in [25] that $\ell_{\infty}(c_{0})$ is isometric to $c_{0,{\sf Fin}^{\omega}}$, and since $\ell_{\infty}(c_{0})$ clearly contains a complemented copy of $c_{0}$, the same holds for $c_{0,{\sf Fin}^{\omega}}$.
+This contrasts with the situation in $\ell_{\infty}$, where every subspace isomorphic or isometric to $c_{0}$ is necessarily not complemented.
+This leads to the following result.
+Proposition 6.8.
+Let $\mathcal{I}$ be an ideal on $\mathbb{N}$. If $c_{0,\mathcal{I}}$ has a complemented copy of $c_{0}$, then $\mathcal{I}$ is not complemented.
+Proof.
+Suppose that $\mathcal{I}$ is complemented, that is, $\ell_{\infty}=c_{0,\mathcal{I}}\oplus W$.
+Since $c_{0,\mathcal{I}}$ has a complemented copy of $c_{0}$, there is a subspace $E$ isomorphic to $c_{0}$ such that $c_{0,\mathcal{I}}=E\oplus Z$.
+Thus, $E$ is complemented in $\ell_{\infty}$, which is impossible by Lindenstrauss’s theorem.
+∎
 Now, we present examples of ideals that are not complemented.
-If $\mathcal{I}$ and $\mathcal{J}$ are ideals, their Fubini product ×$\mathcal{I}\times\mathcal{J}$ is the ideal on $\mathbb{N}\times\mathbb{N}$ defined by
+If $\mathcal{I}$ and $\mathcal{J}$ are ideals, their Fubini product $\mathcal{I}\times\mathcal{J}$ is the ideal on $\mathbb{N}\times\mathbb{N}$ defined by
 $A\in\mathcal{I}\times\mathcal{J}\quad\mbox{if and only if}\quad\{m\in\mathbb{N}\,\colon\,\{n\in\mathbb{N}\,\colon\,(m,n)\in A\}\not\in\mathcal{J}\}\in\mathcal{I}.$
-Regarding the Baire property of the Fubini product, in [11] was shown that ×$\mathcal{I}\times{\sf Fin}$ has the Baire property (hence, it is meager) for any ideal $\mathcal{I}$.
+Regarding the Baire property of the Fubini product, in [11] was shown that $\mathcal{I}\times{\sf Fin}$ has the Baire property (hence, it is meager) for any ideal $\mathcal{I}$.
 On the other hand, they also showed that $\sf Fin$$\times\mathcal{I}$ has the Baire property exactly when $\mathcal{I}$ has it.
@@
     $\mathfrak{ad}({\sf Fin}\times\mathcal{I})=\mathfrak{ad}(\mathcal{I}\times{\sf Fin})=2^{\aleph_{0}}$.
-In particular, ${\sf Fin}\times\mathcal{I}$ and ×$\mathcal{I}\times{\sf Fin}$ are not complemented.
+In particular, $\sf Fin$$\times\mathcal{I}$ and $\mathcal{I}\times{\sf Fin}$ are not complemented.
   2. (2)
@@
 In particular, it follows that $\sf Fin$$\times\mathcal{I}$ is not complemented by Corollary 6.2.
-To show that $\mathfrak{ad}(\mathcal{I}\times{\sf Fin})=2^{\aleph_{0}}$, we know that ×$\mathcal{I}\times{\sf Fin}$ has the Baire property by [11, Proposition 2.6].
+To show that $\mathfrak{ad}(\mathcal{I}\times{\sf Fin})=2^{\aleph_{0}}$, we know that $\mathcal{I}\times{\sf Fin}$ has the Baire property by [11, Proposition 2.6].
 Consequently, it is meager by Theorem 2.1.
 Therefore, we have $\mathfrak{ad}(\mathcal{I}\times{\sf Fin})=2^{\aleph_{0}}$ by Lemma 6.1.
-Now, from [18] it follows that ×$\mathcal{I}\times{\sf Fin}$ is not complemented.
-(2) According to [23, Theorem 5.4], we have the isometry $c_{0,\mathcal{I}^{\omega\perp}}\cong c_{0}(c_{0,\mathcal{I}^{\perp}})$.
+Now, from [20] it follows that $\mathcal{I}\times{\sf Fin}$ is not complemented.
+(2) According to [25, Theorem 5.4], we have the isometry $c_{0,\mathcal{I}^{\omega\perp}}\cong c_{0}(c_{0,\mathcal{I}^{\perp}})$.
 By the main result of [5], the space $c_{0}(c_{0,\mathcal{I}^{\perp}})$ contains a complemented copy of $c_{0}$.
@@
 […]
-The only examples of complemented ideals that we know are the strongly $\omega$-maximal ones.
-The only strict arrows we know are the ones mentioned in the diagram.
-To finish, we present the examples showing that the bottom arrows are strict.
-Example 3.8 shows that the first arrow is strict.
-For the second arrow, an example is given by [15, Lemma 2.2, Corollary 4.1].
-Finally, for the last arrow, we mentioned above ${\sf Fin}\times\mathcal{I}$ is not meager whenever $\mathcal{I}$ is a maximal ideal, but $\mathfrak{ad}({\sf Fin}\times\mathcal{I})=2^{\aleph_{0}}$.
+To conclude, we provide examples showing that all the implications in the diagram are strict.
+Example 3.8 shows that the first implication in the lower row is strict; consequently, the corresponding implication in the upper row is also strict.
+Moreover, there exists a complemented ideal that is not $\omega$-maximal [13, Example 2], which shows that the second implication in the upper row is strict.
+For the second implication in the lower row, an example is given by [16, Lemma 2.2 and Corollary 4.1].
+A counterexample to the last downward implication between the two rows of the diagram is given in [13, Example 1].
+Finally, for the last implication in the lower row, recall that $\sf Fin$$\times\mathcal{I}$ is not meager whenever $\mathcal{I}$ is a maximal ideal, but $\mathfrak{ad}({\sf Fin}\times\mathcal{I})=2^{\aleph_{0}}$.
 Acknowledgments
-We would like to thank Alan Dow for suggesting the Example 3.8. We also thank the partial support given by the VIE-grant 4248 of Universidad Industrial de Santander.
+We would like to thank Alan Dow for suggesting the Example 3.8. We also thank for the partial support given by the VIE-grant 4248 of Universidad Industrial de Santander.